### v2

# Numerical investigation of the interior geometry of semiclassical evaporating spherical charged black holes

Abstract
We developed a numerical code which evolves the semiclassical Einstein’s equation (with the quantum stress-energy contribution added as a source term) for the spherically symmetric metric inside an evaporating semiclassical charged black hole. An analytical approximation for the evolving semiclassical metric was recently developed by Ori and Zilberman (and will be briefly overviewed here). We seek to numerically check the validity of this analytical approximation. The Einstein equations in this case are partial differential equations for the two unknown metric functions which fully describe the spherically symmetric metric. We begin our numerical simulation close to the event horizon with regular initial data specified by a variant of the charged Vaidya metric. We then evolve the metric functions deep into the neighborhood of the inner horizon. We explore the results of running this numerical code in several representative cases. Our numerical simulations confirm the validity of the above mentioned analytical approximation in all these cases.
1 Introduction
Kerr and Reissner–Nordström (RN) geometries, describing rotating and spherically symmetric charged black holes, contain an inner horizon (IH). This IH is a null hypersurface that also serves as a Cauchy horizon (CH) – which is the future boundary where solutions to the field equations, based on initial data provided outside the black hole, fail to be unique. Therefore predictions based on the field equations can be made only up to the Cauchy horizon. The geometry is regular up to and across the IH. There is a unique analytical extension of the geometry beyond the IH, which describes a smooth passage to other external universes [1, 2].
Imagine an observer falling into the black hole. The geometry of the black hole interior will necessarily force the observer to cross the IH (because between the two horizons $r$ is timelike coordinate). What will such an observer encounter? Can the observer cross the IH? In the pure Kerr and RN geometries, there is a smooth passage across the IH, leading to the mentioned other external universes. It turns out, however, that if the black hole is perturbed by classical or quantum matter field, the situation might change, as we will now briefly discuss.
Classical matter fields falling into the black hole may effect the regularity of the CH. As was previously demonstrated, the regular geometry at the IH now becomes a null weak curvature singularity that forms at the Cauchy horizon, both in RN [3, 4, 5, 6, 7, 8] and in Kerr [9, 10, 11, 12] black holes.
Quantum perturbation are also expected to have analogous effects on the geometry. These may be studied using the theoretical framework of semiclassical general relativity, considering quantum fields propagating in classical curved spacetime. In semiclassical forumlation, the stress-energy tensor to be inserted in the Einstein equation is the renormalized expectation value of the stress-energy tensor (RSET) associated with the quantum field (in addition to any classical stress-energy tensor, if such exists–such as electromagnetic stress-energy tensor in the case of RN background). The resulting equation is called the semiclassical Einstein equation.
This equation needs to be handled self-consistently. Note that the curvature of spacetime leads to non-zero RSET of the quantum fields, and this RSET in turn deforms the spacetime, an effect called backreaction.
Solving the semiclassical Einstein equation self-consistently is obviously a challenging task. A natural first step towards tackling this problem is to first compute the RSET for a fixed classical background metric, such as Schwarzchild or RN or Kerr. However, since we are interested in the effect of quantum fields on the IH, we shall not consider here the Schwarzchild case. The near IH RSET, In particular its $T_{uu}$ and $T_{vv}$ component was previously computed for the RN case in Ref. [13] and for the Kerr case in Ref. [14, 15]. For closely related papers see also [16, 17, 18, 19, 20, 21, 22].
For simplicity we restrict our attention here to spherical symmetry, hence we shall focus in this work on the case of spherical charged black hole. We shall investigate the backreaction effect of the RSET in the interior of the black hole, with special attention to the near-IH [^1] Whenever we refer to the evolving backreacted geometry, by “near-IH” we mean the region of very large Eddington-like null coordinates $u$ and $v$. geometry.
Recently an approximate analytical solution was developed by Amos Ori and Noa Zilberman (still unpublished). This approximation is based on the fact that astrophysical black holes have very large masses (compared to the Planck mass) and therefore their evaporation rate is very small. We briefly present this analytical approximation for the semiclassically evolving metric functions below.
The mentioned analytical approximation was developed under a key assumption about the behavior of the RSET in the evolving metric: our assumption is that the asymptotic behavior found near the IH in RN is naturally generalized to the evolving backreaction geometry, in a certain manner that we specify at the end of subsection 3.6 below.
Our main goal in this work is to numerically explore this issue – more specifically, to check the validity of the above mentioned analytical approximation. To this end, we developed a numerical code (published as [23]) which allows us to evolve the semiclassical Einstein equation in this model of spherically symmetric metric that evolves inside the evaporating semiclassical charged black hole. We work in double-null coordinates, using the outgoing null coordinate $u$ and the ingoing null coordinate $v$. The resulting equations are partial differential equations for the two unknown functions $R(u,v)$ and $S(u,v)$ (to be defined later), which fully define the evolving spherically symmetric metric. We begin our numerical simulation close to the EH, and evolve the Einstein’s evolution equations towards the future, until we get deep into the near-IH region. The numerical simulation assumes a specific form for the mass $M(v)$ of the evaporating black hole as a function of the ingoing Eddington coordinate $v$, and that the charge-to-mass ratio $\frac{Q}{M}$ at the EH remains constant upon evaporation – as well as a form for the source terms which incorporate the semiclassical contributions to $T_{uv}$ and $T_{\theta\theta}$, and for the current $J_{u}$  [^2] In practice, we choose a function $Q(u,v)$ which determines the current components $J_{u}$ and $J_{v}$. Our numerical simulation confirms the validity of the mentioned analytical approximation.
Part I Theoretical overview
2 Theoretical background
2.1 Spherically symmetric Einstein equations
Begin with the following spherically-symmetric metric:
$ds^{2}=-2e^{S(u,v)}\frac{L_{0}}{r(u,v)}du\;dv+r^{2}(u,v)d\Omega^{2}$ (2.1)
where $d\Omega^{2}$is defined as:
$d\Omega^{2}=d\theta^{2}+\sin^{2}\theta d\phi^{2}$ (2.2)
We’ll use geometrized unit system where $c=1$ and $G=1$. In this unit system, there is an equivalence between distance, time and mass units. This metric depends on two unknown functions: $S(u,v)$ and $r(u,v)$. $r(u,v)$ has units of length, and so has $L_{0}$. $S(u,v)$ is dimensionless. The coordinates $u$,$v$ have dimensions of length. We’re free to choose $L_{0}$ as we wish. We define
$R(u,v)=r^{2}(u,v)$ (2.3)
Calculating Einstein’s tensor for this metric, using Ricci package in Mathematica we get:
$G_{uu}=\frac{1}{R}\left(R_{,u}S_{,u}-R_{,uu}\right)$ (2.4)
$G_{vv}=\frac{1}{R}\left(R_{,v}S_{,v}-R_{,vv}\right)$ (2.5)
$G_{uv}=\frac{L_{0}e^{S}}{R^{3/2}}+\frac{R_{,uv}}{R}$ (2.6)
$G_{\theta\theta}=-\frac{1}{2L_{0}}e^{-S}\sqrt{R}\left(R_{,uv}+2RS_{,uv}\right)$ (2.7)
The Einstein equation is:
$G_{\mu\nu}=8\pi T_{\mu\nu}$ (2.8)
Those equations lead to the following equations for $R_{,uv}$, $S_{,uv}$:
$R_{,uv}=-\frac{L_{0}e^{S}}{\sqrt{R}}+8\pi RT_{uv}$ (2.9)
$S_{,uv}=\frac{L_{0}e^{S}}{2R^{\frac{3}{2}}}-8\pi\left(\frac{1}{2}T_{uv}+\frac{L_{0}e^{S}}{R^{\frac{3}{2}}}T_{\theta\theta}\right)$ (2.10)
And the following equations for $T_{vv}$, $T_{uu}$:
$8\pi T_{uu}=\frac{1}{R}\left(R_{,u}S_{,u}-R_{,uu}\right)$ (2.11)
$8\pi T_{vv}=\frac{1}{R}\left(R_{,v}S_{,v}-R_{,vv}\right)$ (2.12)
The equations for $R_{,uu}$ and $R_{,vv}$, entailed in Eqs. (2.11,2.12), are the constraint equations, and the equations (2.9,2.10) for $R_{,uv}$$S_{,uv}$ are the evolution equations.
In our model, the stress-energy tensor $T_{\mu\nu}$ is composed of two parts, the electromagnetic contribution $T_{\mu\nu}^{(EM)}$, and the semiclassical part $T_{\mu\nu}^{(SC)}$:
$T_{\mu\nu}=T_{\mu\nu}^{(EM)}+T_{\mu\nu}^{(SC)}$ (2.13)
2.2 Solution to Maxwell’s equations
Assume that the only non-zero component of $F^{\mu\nu}$is $F^{uv}$, as it is required by spherical symmetry. Also assume it is a function of only $u$,$v$ as it is required by spherical symmetry.
The homogeneous Maxwell equation is trivially satisfied:
$F_{\alpha\beta;\gamma}+F_{\beta\gamma;\alpha}+F_{\gamma\alpha;\beta}=0$ (2.14)
It is equivalent to
$F_{\alpha\beta,\gamma}+F_{\beta\gamma,\alpha}+F_{\gamma\alpha,\beta}=0$ (2.15)
It is easy to see the equation is trivially satisfied.
The source-free inhomogeneous equation is:
$0=F_{\quad;\beta}^{\alpha\beta}=\frac{1}{\sqrt{|g|}}\left(\sqrt{|g|}F^{\alpha\beta}\right)_{,\beta}$ (2.16)
Where $g$ is the determinant of the metric, satisfying
$\sqrt{|g|}=r^{2}\sin\theta|g_{uv}|$ (2.17)
Taking the $\alpha=u$ component of the equation:
$\left(r^{2}|g_{uv}|F^{uv}\right)_{,v}=0$ (2.18)
Taking the $\alpha=v$ component of the equation Eq. (2.16) yields an analogous result:
$\left(r^{2}|g_{uv}|F^{uv}\right)_{,u}=0$ (2.19)
From these two equations we get that $r^{2}|g_{uv}|F^{uv}$ is constant, which we call $Q$. $Q$ is the electric charge.
Therefore,
$F^{uv}=\frac{Q}{r^{2}|g_{uv}|}$ (2.20)
We lower the indices to get:
$F_{uv}=g_{uv}^{2}F^{vu}=-\frac{Q}{r^{2}|g_{uv}|}g_{uv}^{2}=-\frac{Q}{r^{2}}|g_{uv}|$ (2.21)
In general, we want to have currents, so we will use non-constant $Q:$
$F_{uv}=-\frac{Q(u,v)}{r^{2}}|g_{uv}|$ (2.22)
$F^{uv}=\frac{Q(u,v)}{r^{2}|g_{uv}|}$ (2.23)
The inhomogeneous Maxwell’s equation with sources is:
$4\pi J^{\alpha}=F_{\quad;\beta}^{\alpha\beta}=\frac{1}{\sqrt{|g|}}\left(\sqrt{|g|}F^{\alpha\beta}\right)_{,\beta}$ (2.24)
Choosing $\alpha=u$:
$4\pi J^{u}=\frac{1}{r^{2}\sin\theta|g_{uv}|}\left(r^{2}\sin\theta|g_{uv}|\frac{Q(u,v)}{r^{2}|g_{uv}|}\right)_{,v}=\frac{Q_{,v}}{r^{2}|g_{uv}|}$ (2.25)
Choosing $\alpha=v$ and doing a similar calculation:
$4\pi J^{v}=-\frac{Q_{,u}}{r^{2}|g_{uv}|}$ (2.26)
Lowering the indices, we get:
$4\pi J_{u}=g_{uv}\frac{-Q_{,u}}{r^{2}|g_{uv}|}=\frac{Q_{,u}}{r^{2}}$ (2.27)
$4\pi J_{v}=-\frac{Q_{,v}}{r^{2}}$ (2.28)
2.3 The electromagnetic stress-energy tensor
We can now calculate the electromagnetic stress-energy tensor for this given electromagnetic field tensor:
$T_{\mu\nu}^{(EM)}=\frac{1}{4\pi}\left(F_{\mu\lambda}F_{\nu}^{\;\lambda}-\frac{1}{4}g_{\mu\nu}F_{\lambda\kappa}F^{\lambda\kappa}\right)$ (2.29)
Noting that:
$F_{\lambda\kappa}F^{\lambda\kappa}=2F_{uv}F^{uv}=-2\frac{Q}{r^{2}|g_{uv}|}\frac{Q}{r^{2}}|g_{uv}|=-\frac{2Q^{2}}{r^{4}}$ (2.30)
We obtain
$T_{uv}^{(EM)}$ $=\frac{1}{4\pi}\left(F_{uv}g^{uv}F_{vu}-\frac{1}{4}g_{uv}F_{\lambda\kappa}F^{\lambda\kappa}\right)$ (2.31)
$=\frac{1}{4\pi}\left(-\frac{Q^{2}}{r^{4}}|g_{uv}|^{2}g^{uv}-\frac{1}{4}g_{uv}F_{\lambda\kappa}F^{\lambda\kappa}\right)$ (2.32)
$=\frac{1}{4\pi}\left(-\frac{Q^{2}}{r^{4}}g_{uv}+\frac{1}{4}g_{uv}\frac{2Q^{2}}{r^{4}}\right)=-\frac{1}{8\pi}\frac{Q^{2}}{r^{4}}g_{uv}$ (2.33)
Recalling that
$g_{uv}=-\frac{e^{S}L_{0}}{r}$ (2.34)
We get:
$T_{uv}^{(EM)}=\frac{1}{8\pi}\frac{Q^{2}e^{S}L_{0}}{r^{5}}$ (2.35)
Calculating $T_{\theta\theta}^{(EM)}$, we note that the first term $F_{\mu\lambda}F_{\nu}^{\;\lambda}$ trivially vanishes in this case, getting:
$T_{\theta\theta}^{(EM)}=\frac{1}{4\pi}\left[-\frac{1}{4}g_{\theta\theta}\left(-\frac{2Q^{2}}{r^{4}}\right)\right]=\frac{1}{8\pi}\frac{Q^{2}}{r^{2}}$ (2.36)
Because of spherical symmetry, $T_{\phi\phi}=T_{\theta\theta}\sin^{2}\theta$, for both the electromagnetic stress-energy tensor and the semi-classical stress energy tensor. Note that $T_{vv}^{(EM)}=T_{uu}^{(EM)}=0$, therefore $T_{uu}=T_{uu}^{(SC)}$ and $T_{vv}=T_{vv}^{(SC)}$. We will hereafter refer to $T_{uu}$ and $T_{vv}$ without the “(SC)(SC)” mark.
Using Einstein’s equation $G_{\mu\nu}=8\pi T_{\mu\nu}$, and assuming the only stress-energy tensor is the electromagnetic one, the evolution equations in the presence of charge are, based on equations Eq. (2.9,2.10)
$R_{,uv}=-\frac{e^{S}L_{0}}{\sqrt{R}}+e^{S}L_{0}\frac{Q^{2}}{R^{\frac{3}{2}}}=e^{S}L_{0}F_{1}(R)$ (2.37)
$S_{,uv}=\frac{e^{S}L_{0}}{2R^{\frac{3}{2}}}-\frac{3}{2}e^{S}L_{0}\frac{Q^{2}}{R^{\frac{5}{2}}}=e^{S}L_{0}F_{2}(R)$ (2.38)
where
$F_{1}(R)=-\frac{1}{\sqrt{R}}+\frac{Q^{2}}{R^{\frac{3}{2}}}$ (2.39)
$F_{2}(R)=\frac{1}{2R^{\frac{3}{2}}}-\frac{3}{2}\frac{Q^{2}}{R^{\frac{5}{2}}}$ (2.40)
Notice that:
$F_{2}(R)=\frac{d}{dR}F_{1}(R)$ (2.41)
In view of this property, the PDE system (2.37,2.38) belongs to the general class of PDE systems investigated in Ref. [24].
Note that we here assumed that the only stress-energy tensor is the electromagnetic one. We will later break this assumption, by taking into account the non-zero $T_{\mu\nu}^{(SC)}$. This will lead to new source terms in the evolution equations, see subsection 2.8.
Taking the constraint equation Eq. (2.12) we get:
$\widetilde{T}_{vv}\equiv 8\pi r^{2}T_{vv}=R_{,v}S_{,v}-R_{,vv}$ (2.42)
and similarly for $T_{uu}:$
$\widetilde{T}_{uu}\equiv 8\pi r^{2}T_{uu}=R_{,u}S_{,u}-R_{,uu}$ (2.43)
Here we defined the two quantities $\widetilde{T}_{vv},\widetilde{T}_{uu}.$
The quantity $\widetilde{T}_{uu}$ satisfies a conservation equation :
$\left(\widetilde{T}_{uu}\right)_{,v}=0$ (2.44)
To prove this, we differentiate Eq. (2.43) with respect to $v$, and use the evolution equations Eq. (2.9,2.10):
$\left(\widetilde{T}_{uu}\right)_{,v}$ $=R_{,uv}S_{,u}+R_{,u}S_{,uv}-\left(R_{,uv}\right)_{,u}$ (2.45)
$=e^{S}L_{0}F_{1}(R)S_{,u}+e^{S}L_{0}F_{2}(R)R_{,u}-\left[e^{S}L_{0}F_{1}(R)\right]_{,u}$ (2.46)
$=e^{S}L_{0}F_{1}(R)S_{,u}+e^{S}L_{0}F_{2}(R)R_{,u}-e^{S}L_{0}F_{1}(R)S_{,u}-e^{S}L_{0}\frac{dF_{1}(R)}{dR}R_{,u}=0$ (2.47)
In the last equation we used the relation Eq. (2.41) between $F_{1}$ and $F_{2}$. Note that we assumed $Q$ is constant at this stage. When $Q$ changes the situation changes and we address it in section 3.6.
In the same manner, we can derive the analogous conservation equation for $\widetilde{T}_{vv}$ :
$\left(\widetilde{T}_{vv}\right)_{,u}=0$ (2.48)
Note, however, that when we add the semiclassical source terms Eqs. (2.44, 2.48) will not be valid.
2.4 Reissner–Nordström metric
The Reissner–Nordström metric is:
$ds^{2}=-\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)dt^{2}+\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)^{-1}dr^{2}+r^{2}d\Omega^{2}$ (2.49)
$r_{+},r_{-},\kappa_{+},\kappa_{-}$ are defined as follows:
$r_{+}=M+\sqrt{M^{2}-Q^{2}}$ (2.50)
$r_{-}=M-\sqrt{M^{2}-Q^{2}}$ (2.51)
$\kappa_{+}=\frac{r_{+}-r_{-}}{2r_{+}^{2}}=\frac{\sqrt{M^{2}-Q^{2}}}{\left(\sqrt{M^{2}-Q^{2}}+M\right)^{2}}$ (2.52)
$\kappa_{-}$ $=\frac{r_{+}-r_{-}}{2r_{-}^{2}}=\frac{\sqrt{M^{2}-Q^{2}}}{\left(M-\sqrt{M^{2}-Q^{2}}\right)^{2}}$ (2.53)
$=\frac{1}{M}\frac{\sqrt{1-\left(\frac{Q}{M}\right)^{2}}}{\left(1-\sqrt{1-\left(\frac{Q}{M}\right)^{2}}\right)^{2}}$ (2.54)
Note that:
$f\equiv\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)=\frac{\left(r-r_{-}\right)\left(r-r_{+}\right)}{r^{2}}$ (2.55)
where we defined $f(r)$. The horizons are the surfaces where $g_{uv}=0$. They are $r=r_{+}$at the event horizon (EH) and $r=r_{-}$ at the inner horizon (IH).
2.5 Ingoing Eddington coordinates
We wish to move to different coordinates. For this purpose we define $r_{*}$:
$r_{*}=r-M+\frac{1}{2\kappa_{+}}\log\left|\frac{r-r_{+}}{M-r_{+}}\right|-\frac{1}{2\kappa_{-}}\log\left|\frac{r-r_{-}}{M-r_{-}}\right|$ (2.56)
Note that:
$\frac{\partial r_{*}}{\partial r}=1+\frac{1}{2\kappa_{+}\left(r-r_{+}\right)}-\frac{1}{2\kappa_{-}\left(r-r_{-}\right)}=\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)^{-1}$ (2.57)
and the constant of integration was chosen such that $r_{*}(r=M)=0$. We move from $(t,r)$ to $(v,r)$ coordinates inside the BH defined by
$v=r_{*}(r)+t$ (2.58)
The inverse coordinate transformation is:
$t=v-r_{*}(r)$ (2.59)
And its partial derivatives are:
$\frac{\partial t}{\partial v}=1,\quad\frac{\partial r}{\partial v}=0$ (2.60)
$\frac{\partial t}{\partial r}=-\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)^{-1},\quad\frac{\partial r}{\partial r}=1$ (2.61)
The metric changes according to:
$g_{vv}=\frac{\partial t}{\partial v}\frac{\partial t}{\partial v}g_{tt}+\frac{\partial r}{\partial v}\frac{\partial r}{\partial v}g_{rr}=-\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)$ (2.62)
$g_{rr}$ $=\frac{\partial t}{\partial r}\frac{\partial t}{\partial r}g_{tt}+\frac{\partial r}{\partial r}\frac{\partial r}{\partial r}g_{rr}$ (2.63)
$=-\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)^{-2}\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)$ (2.64)
$+\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)^{-1}=0$ (2.65)
And last metric component:
$g_{rv}$ $=\frac{\partial t}{\partial r}\frac{\partial t}{\partial v}g_{tt}+\frac{\partial r}{\partial r}\frac{\partial r}{\partial v}g_{rr}$ (2.66)
$=\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)^{-1}\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)=1$ (2.67)
Thus the RN geometry in ingoing Eddington coordinates $(v,r)$ takes the form:
$ds^{2}=-\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)dv^{2}+2dr\;dv+r^{2}d\Omega^{2}$ (2.68)
2.6 RN in double null Eddington coordinates
We wish to move to double-null coordinates inside the black hole. First, we’ll do a coordinate transformation from $(t,r)$ to $(t,r_{*})$. Recalling that:
$\frac{\partial r_{*}}{\partial r}=\left(1-\frac{2M}{r(r_{*})}+\frac{Q^{2}}{r^{2}(r_{*})}\right)^{-1}$ (2.69)
we get:
$g_{r_{*}r_{*}}=g_{rr}\frac{\partial r}{\partial r_{*}}\frac{\partial r}{\partial r_{*}}=\left(1-\frac{2M}{r(r_{*})}+\frac{Q^{2}}{r^{2}(r_{*})}\right)$ (2.70)
The metric in $(t,r_{*})$ coordinates is thus:
$ds^{2}=\left(1-\frac{2M}{r(r_{*})}+\frac{Q^{2}}{r^{2}(r_{*})}\right)\left(-dt^{2}+dr_{*}^{2}\right)+r^{2}(r_{*})d\Omega^{2}$ (2.71)
Now we move to double null Eddington coordinates $(u,v)$:
$v=r_{*}+t$ (2.72)
$u=r_{*}-t$ (2.73)
From these we get the inverse transformation:
$r_{*}=\frac{u+v}{2}$ (2.74)
$t=\frac{v-u}{2}$ (2.75)
We get the following relation for the differentials:
$dv=dr_{*}+dt$ (2.76)
$du=dr_{*}-dt$ (2.77)
Multiplying these together we get:
$du\cdot dv=dr_{*}^{2}-dt^{2}$ (2.78)
Substituting this in the metric Eq. (2.71), the Reissner–Nordström geometry in double null coordinates is found to take the form:
$ds^{2}=\left(1-\frac{2M}{r(u,v)}+\frac{Q^{2}}{r^{2}(u,v)}\right)du\;dv+r^{2}(u,v)d\Omega^{2}$ (2.79)
The metric function $g_{uv}$ is equal to:
$g_{uv}=\frac{1}{2}\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)$ (2.80)
The function $r(u,v)$ is given by inverting the function $r_{*}(r)$:
$\frac{u+v}{2}=r_{*}=r-M+\frac{1}{2\kappa_{+}}\log\left|\frac{r-r_{+}}{M-r_{+}}\right|-\frac{1}{2\kappa_{-}}\log\left|\frac{r-r_{-}}{M-r_{-}}\right|$ (2.81)
We wish to find an expression for the derivatives of the two functions $R(u,v)$ and $S(u,v)$ in RN near the IH. Around $r=r_{-},$we get:
$\frac{u+v}{2}=r_{*}=const-\frac{1}{2\kappa_{-}}\log\left|r-r_{-}\right|$ (2.82)
Rearranging the terms, we get:
$r-r_{-}=const\cdot e^{-2\kappa_{-}r_{*}}=const\cdot e^{-\kappa_{-}\left(u+v\right)}$ (2.83)
The metric component $g_{uv}$ (Eq. (2.80)) around $r=r_{-}$ is (recalling also Eq. (2.1,2.55)):
$2\frac{e^{S}L_{0}}{r}$ $=-2g_{uv}$ (2.84)
$=-\frac{\left(r-r_{-}\right)\left(r-r_{+}\right)}{r^{2}}\approx-\frac{\left(r-r_{-}\right)\left(r_{-}-r_{+}\right)}{r_{-}^{2}}$ (2.85)
$=2\kappa_{-}\left(r-r_{-}\right)$ (2.86)
Combining these equations we get:
$const\cdot 2\kappa_{-}e^{-\kappa_{-}\left(u+v\right)}=2\kappa_{-}\left(r-r_{-}\right)\approx 2\frac{e^{s}L_{0}}{r_{-}}$ (2.87)
Taking the log of this equation,
$S=-\kappa_{-}\left(u+v\right)+const$ (2.88)
So:
$S_{,u}=S_{,v}=-\kappa_{-}$ (2.89)
near the IH. Taking the derivative of Eq. (2.83) with respect to $r_{*},$we get near the IH:
$r_{,r_{*}}=const\cdot e^{-2\kappa_{-}r_{*}}$ (2.90)
Using this, we can calculate the derivatives of $R$ around $r=r_{-}:$
$R_{,u}=R_{,r}r_{,r_{*}}r_{*,u}=2r\cdot const\cdot e^{-2\kappa_{-}r_{*}}\frac{1}{2}=const\cdot e^{-2\kappa_{-}r_{*}}\rightarrow 0$ (2.91)
as $r_{*}\rightarrow\infty$, and similarly for $R_{,v}$:
$R_{,v}=R_{,r}r_{,r_{*}}r_{*,v}=const\cdot e^{-2\kappa_{-}r_{*}}\rightarrow 0$ (2.92)
In a similar manner, around the EH, at $r\rightarrow r_{+}$ from below, we get:
$r_{*}=const+\frac{1}{2\kappa_{+}}\log\left|r-r_{+}\right|$ (2.93)
Rearranging the terms, we get:
$r-r_{+}=const\cdot e^{2\kappa_{+}r_{*}}=const\cdot e^{\kappa_{+}\left(u+v\right)}$ (2.94)
The metric component $g_{uv}$ around $r=r_{+}$is :
$2\frac{e^{S}L_{0}}{r}$ $=-2g_{uv}$ (2.95)
$=-\frac{\left(r-r_{+}\right)\left(r-r_{-}\right)}{r^{2}}\approx-\frac{\left(r-r_{+}\right)\left(r_{+}-r_{-}\right)}{r_{-}^{2}}$ (2.96)
$=-2\kappa_{+}\left(r-r_{+}\right)$ (2.97)
Combining these equations we get:
$const\cdot 2\kappa_{+}e^{\kappa_{+}\left(u+v\right)}=2\kappa_{+}\left(r-r_{+}\right)\approx 2g_{uv}$ (2.98)
2.7 Vaidya metric
Our solution near the EH will be rooted by the ingoing charged Vaidya metric.[25]. The Vaidya metric is:
$ds^{2}=-\left(1-\frac{2M(v)}{r}+\frac{Q(v)^{2}}{r^{2}}\right)dv^{2}+2dvdr+r^{2}d\Omega^{2}$ (2.99)
Note that this metric, when taking $M(v)=const$ and $Q(v)=const$, is exactly equal to Eq. (2.68) ingoing Eddington metric.
The $G_{vv}$ component of the Einstein tensor of the Vaidya metric is:
$G_{vv}=\frac{1}{r^{2}}\left(2M_{,v}-2\frac{1}{r}QQ_{,v}\right)+\left(1-\frac{2M(v)}{r}+\frac{Q(v)^{2}}{r^{2}}\right)\frac{Q^{2}}{r^{4}}$ (2.100)
The $G_{rv}$ component of the Einstein tensor of the Vaidya metric is:
$G_{rv}=-\frac{Q^{2}}{r^{4}}$ (2.101)
We wish to know $G_{vv}$ in double null coordinates $(v,u)$, instead of $v,r$. Double null coordinates are characterized by $g_{uu}=g_{vv}=0$.
$0$ $=g_{vv}=\frac{\partial v}{\partial v}\frac{\partial v}{\partial v}g_{vv}+2\frac{\partial r}{\partial v}\frac{\partial v}{\partial v}g_{rv}$ (2.102)
$=-\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)+2\frac{\partial r}{\partial v}$ (2.103)
Transformations into double null coordinates are therefore characterized by:
$\frac{\partial r}{\partial v}=\frac{1}{2}\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)$ (2.104)
Therefore, $G_{vv}$ is equal to
$G_{vv}$ $=\frac{\partial v}{\partial v}\frac{\partial v}{\partial v}G_{vv}+2\frac{\partial r}{\partial v}\frac{\partial v}{\partial v}G_{rv}$ (2.105)
$=\frac{1}{r^{2}}\left(2M_{,v}-2\frac{1}{r}QQ_{,v}\right)$ (2.106)
Therefore, in double null coordinates,
$\widetilde{T}_{vv}=r^{2}G_{vv}=2M_{,v}-2\frac{1}{r}QQ_{,v}$ (2.107)
The inverse metric tensor is:
$g^{rr}=1-\frac{2M(v)}{r}+\frac{Q(v)^{2}}{r^{2}}$ (2.108)
$g^{rv}=1$ (2.109)
$g^{vv}=0$ (2.110)
We raise the indices to get $G^{vv}$:
$G^{vv}=g^{v\alpha}g^{v\beta}G_{\alpha\beta}=0$ (2.111)
$G^{vv}$ is the same in double null coordinates:
$G^{vv}=\frac{\partial v}{\partial v}\frac{\partial v}{\partial v}G^{vv}=0$ (2.112)
We lower the indices to get $G_{UU}$:
$G_{uu}=g_{uv}g_{uv}G^{vv}=0$ (2.113)
Therefore,
$\widetilde{T}_{uu}=0$ (2.114)
2.8 The source terms in the evolution equations
As mentioned in Eqs. (2.44,2.48), in the absence of $T_{uv}^{(SC)},T_{\theta\theta}^{(SC)}$ (namely in double null fluid model) and with constant charge ($Q=const$), $\widetilde{T}_{uu,v}=\widetilde{T}_{vv,u}=0$ – that is, both $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ are individually conserved. Since our initial values are taken from the Vaidya geometry near the EH, and since $\widetilde{T}_{uu}$ vanishes there (as is evident from regularity conditions), this implies that $\widetilde{T}_{uu}=0$ throughout. We also have $\widetilde{T}_{vv}=2M_{,v}$ , see Eq. (2.107).
When the charge is changing and there is an electric current – which is the general case we are actually considering in this work – if the current is only $J_{v}$ while $J_{u}=0$, then $\widetilde{T}_{vv}=2M_{,v}-2\frac{1}{r}QQ_{,v}$ but $\widetilde{T}_{uu}=0$ still holds.
As a matter of fact, however, the analysis of semi-classical fluxes at the inner horizon [13] (for uncharged scalar field, yielding $J_{\mu}=0$) showed that generically $T_{uu}\neq 0$ there. From the above discussion, it follows that this is only possible because at the intermediate region between $r_{+}$ and $r_{-}$ there are non-zero $T_{uv}^{(SC)},T_{\theta\theta}^{(SC)}$. Thus, to allow us modeling the fact that $T_{uu}\neq 0$ at the inner horizon, we must add such non-zero $T_{uv}^{(SC)},T_{\theta\theta}^{(SC)}$ in the region between the EH and the IH.
The evolution equations in the presence of non-vanishing $T_{uv}^{(SC)},T_{\theta\theta}^{(SC)}$ take the form:
$R_{,uv}=e^{S}L_{0}F_{1}(R)+Z_{R}$ (2.115)
$S_{,uv}=e^{S}L_{0}F_{2}(R)+Z_{S}$ (2.116)
where $Z_{R}$$Z_{S}$ are the source terms in the evolution equation:
$Z_{R}=8\pi RT_{uv}^{(SC)}$ (2.117)
$Z_{S}=-8\pi\left(\frac{1}{2}T_{uv}^{(SC)}+\frac{L_{0}e^{S}}{R^{\frac{3}{2}}}T_{\theta\theta}^{(SC)}\right)$ (2.118)
Note that the constraint equations remain unchanged, they do not depend at all on $T_{uv}^{(SC)}$ and $T_{\theta\theta}^{(SC)}$.
Obviously, in order to run a numerical simulation, we have to make a concrete choice of $T_{uv}^{(SC)}$ and $T_{\theta\theta}^{(SC)}$ (as functions of $u$ and $v$). Unfortunately, these two semiclassical functions $T_{uv}^{(SC)},T_{\theta\theta}^{(SC)}$ have not been computed yet inside black holes. Naively one might view this lack of knowledge of $T_{uv}^{(SC)},T_{\theta\theta}^{(SC)}$ as a crucial obstacle. This is not the case, however, as we now explain:
The purpose of our numerical simulation is to test the validity of an analytical approximation for the evolving semiclassical metric (particularly near the IH), which will be described later (see Sec. 4). This analytical approximation only depends on a set of four main semiclassical input functions: $M(v)$ (closely related to $\widetilde{T}_{vv}^{(+)}(v)$) and $Q^{v}(v)$ (closely related to $J_{v}^{(+)}(v)\propto Q_{,v}^{v}$) at the event horizon – and $\widetilde{T}_{uu}^{(-)}(v),$$\text{$\widetilde{T}_{vv}^{(-)}$}(v)$ at the inner horizon, both defined with respect to a background RN metric with parameters $M(v)$,$Q(v)$ . This analytical approximation does not depend on $T_{uv}^{(SC)}$ and $T_{\theta\theta}^{(SC)}$. [The actual values of $\widetilde{T}_{uu}^{(-)},$$\widetilde{T}_{vv}^{(-)}$ are affected by the presence of $T_{uv}^{(SC)},T_{\theta\theta}^{(SC)}$ between the two horizon; but nevertheless, once $\widetilde{T}_{uu}^{(-)},\text{$\widetilde{T}_{vv}^{(-)}$}$ are known (and indeed they have been computed, e.g. in Ref. [13]), we no longer need $T_{uv}^{(SC)},T_{\theta\theta}^{(SC)}$ for the mentioned analytical approximation.]
The way we proceed in our numerical investigation is therefore the following: We arbitrarily guess the two semiclassical quantities $T_{uv}^{(SC)},T_{\theta\theta}^{(SC)}$ (as functions of $u,v$), and then numerically evolve the evolution equations (starting from initial data prescribed near the EH). The choice of these two RSET components is given in subsection 3.6. (Overall we numerically test 7 different configurations for these RSET components – all of them are of order comparable to the fluxes associated with the evaporation process.) In this process we obtain the evolving unknown functions $R(u,v)$ and $S(u,v)$ – and thereby also the two near-IH flux functions $\widetilde{T}_{uu}^{(-)},$$\widetilde{T}_{vv}^{(-)}$ from the constraint equations Eqs. (2.42, 2.43) . We can then substitute these two flux functions in the analytical approximation, and compare it to the actual numerically-evolving functions $R(u,v)$ and $S(u,v)$. This method allows us to numerically test the analytical approximation for any (quite arbitrary) choice of $T_{uv}^{(SC)}$ and $T_{\theta\theta}^{(SC)}$.
3 Inputs from semiclassical theory
3.1 Hawking radiation
We start by discussing Schwarzschild black hole. According to Hawking [26], in the presence of quantum fields, the black hole radiates away particles and will eventually evaporate. According to Hawking, the black hole has Hawking temperature of $T_{H}$ (we’re using units where Boltzmann constant $k_{B}=1$):
$T_{H}=\frac{\hbar\kappa_{+}}{2\pi}=\frac{\hbar}{8\pi M}$ (3.1)
The most naive expression for the energy radiation rate is:
$L=\sigma AT_{H}^{4}$ (3.2)
where $\sigma=\frac{\pi^{2}}{60\hbar^{3}}$ is the Stefan-Boltzmann constant, and $A=4\pi r_{+}^{2}=16\pi M^{2}$ is the area of the sphere of the EH. In Schwarzschild, this luminosity is equal to:
$L=\frac{\pi^{2}}{60\hbar^{3}}16\pi M^{2}\frac{\hbar^{4}}{8^{4}\pi^{4}M^{4}}=\frac{\hbar}{15360\pi M^{2}}$ (3.3)
However, black holes aren’t perfect black bodies, and some waves get reflected by the black hole. The actual luminosity is therefore equal to:
$L=\frac{\hbar}{M^{2}}P$ (3.4)
where $P$ is some dimensionless constant, which needs to be computed numerically. According to Page [27], the constant is equal to $P=3.366\times 10^{-5}$ for photon emissions, and $P=0.3845\times 10^{-5}$ for graviton emissions, and according to Elster [28] $P=7.44\times 10^{-5}$ for scalar field. There are two degrees of freedom for the photon and graviton while there is a single degree of freedom for the scalar.
In the case of RN black hole, the luminosity depends also on the charge, and it is of the form:
$L=\frac{\hbar}{M^{2}}H_{M}\left(\frac{Q}{M}\right)$ (3.5)
for some dimensionless function $H_{M}\left(\frac{Q}{M}\right)$ of the variable $\frac{Q}{M}$ (that may be determined numerically).
The luminosity determines the rate of change of the mass $M_{,t}$, as observed by a distant observer:
$M_{,t}^{(\infty)}=-L$ (3.6)
At the EH, the rate of change of the mass in the Eddington coordinate $v$ is the same as the rate of change of mass seen by a distant observer:
$M_{,v}^{(+)}=M_{,t}^{(\infty)}=-L=-\frac{\hbar}{M^{2}}H_{M}\left(\frac{Q}{M}\right)$ (3.7)
In this entire work we will assume, for simplicity, that $\frac{Q}{M}$ remains constant upon evaporation of the black hole. We will discuss this assumption later. We can then solve this differential equation for $M(v)$ and get:
$M(v)=\left(M_{0}^{3}-3v\hbar H_{M}\left(\frac{Q}{M}\right)\right)^{\frac{1}{3}}$ (3.8)
where $M_{0}$ is the initial mass at $v=0$. We’ll define the dimensionless parameter $\varepsilon_{0}\equiv 3\hbar M_{0}^{-2}H_{M}\left(\frac{Q}{M}\right)$ and obtain:
$M(v)=\left(M_{0}^{3}-M_{0}^{2}\varepsilon_{0}v\right)^{\frac{1}{3}}$ (3.9)
Therefore the evaporation rate is:
$M_{,v}=-\frac{1}{3}M_{0}^{2}\varepsilon_{0}\left(M_{0}^{3}-M_{0}^{2}\varepsilon_{0}v\right)^{-\frac{2}{3}}$ (3.10)
The evaporation rate at the beginning is:
$M_{,v}(v=0)=-\frac{1}{3}\varepsilon_{0}$ (3.11)
The evaporation time $T_{ev}$ from the initial moment (when $M=M_{0}$) until the moment when $M(v)=0$ is:
$T_{ev}=\frac{M_{0}^{3}}{3\hbar H_{M}\left(\frac{Q}{M}\right)}$ (3.12)
A convenient dimensionless quantity that characterizes the initial mass size is $\frac{T_{ev}}{M_{0}}$:
$\frac{T_{ev}}{M_{0}}=\frac{M_{0}^{2}}{3\hbar H_{M}\left(\frac{Q}{M}\right)}=\frac{1}{\varepsilon_{0}}$ (3.13)
3.2 Choice of initial mass parameter
Since we wish to cover significant evaporation in our numerical simulation (we typically manage to evaporate about 53% of the initial mass and charge), our numerical grid must cover a grid length of order $\frac{T_{ev}}{M_{0}}\left(\frac{du}{M_{0}}\right)^{-1}$points. ($du$ is the size of a single grid step, and we usually have $\frac{du}{M_{0}}=\frac{1}{100}$.) For astrophysical black holes of 10 solar masses, $\frac{T_{ev}}{M_{0}}\approx 10^{82}$, which is impossible to implement numerically. We can implement numerically $\frac{T_{ev}}{M_{0}}=10^{3}$. This value definitely does not correspond to a realistic astrophysical black hole, yet this value is still $\gg 1$. We’ll therefore choose our $M_{0}$ such that $\frac{T_{ev}}{M_{0}}=10^{3}$. The corresponding value of $\varepsilon_{0}$ is $\frac{1}{1000}$.
3.3 Choice of units
Recall that as mentioned above, we’re using general relativistic geometrized unit system $c=G=1$, which leaves us with a single free unit of mass scale. We can choose this mass unit as we wish, so we’ll choose it to be our initial mass $M_{0}$, that is, $M_{0}=1$. In other words, our mass unit is relative to the initial mass of the black hole.
Recall the parameter $L_{0}$ from Eq. (2.1), which can be chosen freely. We choose $L_{0}=M_{0}$. In our choice of units, we therefore get $L_{0}=1$.
3.4 Basic scaling of semiclassical $T_{\alpha}^{\;\beta}$
Classical general relativity is scale-invariant. In particular, every Schwarzschild black hole is similar to every other Schwarzschild black hole via a mapping which keeps $\frac{r}{M}$ constant. Every physical quantity whose dimensions are $l]^{n}$ scales under this mapping as $M^{n}$. In the case of RN, all black holes with the same $\frac{Q}{M}$ are also similar in that sense.
In the semiclassical theory, quantities may also depend on $\hbar$. In geometrized units ($c=G=1$) $\hbar$ has dimensions of length squared, so $\frac{\hbar}{M^{2}}$ is dimensionless. The semiclassical $T_{\alpha}^{\;\beta}$ has the form [^3] We here refer to the expectation value of the renormalized stress-energy tensor:
$T_{\alpha}^{\;\beta(SC)}=\frac{\hbar}{M^{2}}\hat{T}_{\alpha}^{\;\beta}$ (3.14)
where $\hat{T}_{\alpha}^{\;\beta}$ is independent of $\hbar$, and therefore scales like a classical quantity with the same dimensions as this quantity $\hat{T}_{\alpha}^{\;\beta}$ has. But $\frac{\hbar}{M^{2}}$ is dimensionless and therefore $\hat{T}_{\alpha}^{\;\beta}]=[T_{\alpha}^{\;\beta(SC)}]$. $T_{\alpha}^{\;\beta(SC)}$ (like any kind of $T_{\alpha}^{\;\beta}$) has units of energy density, that is, it has dimensions of $T_{\alpha}^{\;\beta(SC)}]=[l]^{-2}$. Overall, we obtain that $\hat{T}_{\alpha}^{\;\beta}$ scales like $M^{-2}$. We therefore conclude that $T_{\alpha}^{\;\beta(SC)}$ scales as:
$T_{\alpha}^{\;\beta(SC)}=\frac{\hbar}{M^{4}}H_{T_{\alpha}^{\;\beta}}\left(\frac{r}{M},\frac{Q}{M}\right)$ (3.15)
where $H_{T_{\alpha}^{\;\beta}}\left(\frac{r}{M},\frac{Q}{M}\right)$ is some dimensionless function (free of $\hbar$) that only depends on its two dimensionless arguments. We’ll denote further dimensionless functions of this sort by $H_{...}$, where the “...” denotes various possible subscripts which distinguish these functions from each other.
In particular, the scaling of $T_{vv}$ is:
$T_{vv}\equiv T_{vv}^{(SC)}=\frac{\hbar}{M^{4}}H_{T_{vv}}\left(\frac{r}{M},\frac{Q}{M}\right)$ (3.16)
where, again, $H_{T_{vv}}\left(\frac{r}{M},\frac{Q}{M}\right)$ is some dimensionless function (free of $\hbar$) of its two dimensionless arguments.
In Eq. (2.42) we defined the quantity $\widetilde{T}_{vv}\equiv 8\pi r^{2}T_{vv}$. Correspondingly we define $\widetilde{T}_{vv}^{(+)}$ , $\widetilde{T}_{vv}^{(-)}$ , which are $\widetilde{T}_{vv}$ at the two horizons, EH and IH respectively. They scale according to:
$\widetilde{T}_{vv}^{(+)}=8\pi r_{+}^{2}\frac{\hbar}{M^{4}}H_{T_{vv}}\left(\frac{r_{+}}{M},\frac{Q}{M}\right)\equiv\frac{\hbar}{M^{2}}H_{T_{vv}}^{(+)}\left(\frac{Q}{M}\right)$ (3.17)
$\widetilde{T}_{vv}^{(-)}=8\pi r_{-}^{2}\frac{\hbar}{M^{4}}H_{T_{vv}}\left(\frac{r_{-}}{M},\frac{Q}{M}\right)\equiv\frac{\hbar}{M^{2}}H_{T_{vv}}^{(-)}\left(\frac{Q}{M}\right)$ (3.18)
3.5 The charge function $Q(u,v)$
We would like the charge of the black hole to evaporate as well as the mass, so that the mass could evaporate completely. (If the charge does not evaporate, the mass will stop evaporating when the black hole approaches extremality. We are interested in the case of significantly evaporating mass.) If there is a charged quantum field, then $Q$ will indeed evaporate, and we will have electric currents in spacetime. We will therefore assume that there is such a charged quantum field. Hence $Q$ is no longer a constant parameter, but a function $Q(u,v)$ of $u$ and $v$. At the horizon, we have
$Q=Q^{v}(v)$ (3.19)
and in the whole spacetime we have $Q(u,v)$, which for convenience we express as:
$Q(u,v)=Q^{v}(v)+\delta Q(u,v)$ (3.20)
Note that from Eq. (2.27) if $J_{u}=0$ throughout, then $Q_{,u}=0$, and hence $\delta Q(u,v)=0$ because $\delta Q$ vanishes at the EH.
We need to choose the charge evaporation function $Q^{v}(v)$. To simplify the black hole evaporation model, we choose this function such that:
$\frac{Q^{v}(v)}{M(v)}=const$ (3.21)
Note that we choose $0<\frac{Q^{v}}{M}<1$. The derivation of the analytical approximation below in section 4 is not valid in the extremal case because this approximation is based on the assumption $\kappa_{-}>0$.
To what extent is this choice realistic? Let us consider the (arguably hypothetical  [^4] This case is not regarded as realistic because the plasma that typically surrounds black holes would discharge them very quickly.) case of an astrophysically large black hole which is significantly charged initially. Such a black hole would lose charge due to the Schwinger pair production effect, where the strong electric field near the EH causes electron-positron pairs to be created from the vacuum: one of the charged particles would fall into the black hole, reducing its charge, and the other would escape to infinity. This effect, if possible, would quickly discharge the black hole, especially since the charge-to-mass ratio for electrons is $\approx 2\cdot 10^{21}$. A detailed analysis of schwinger pair production can be found in Ref [29]. Schwinger pair production is effective only if the electric field at the event horizon exceeds a critical value $E_{cr}$. The electric field at the EH near extremality is of order $E_{eh}=\frac{Q}{r_{+}^{2}}\approx\frac{Q}{M^{2}}=\frac{Q}{M}\frac{1}{M}$. Therefore, if $\frac{Q}{M}\frac{1}{M}>E_{cr}$, Schwinger pair production will quickly discharge the black hole until $\frac{Q}{M}\frac{1}{M}=E_{cr}$. This means there is a critical mass $M_{cr}=\frac{1}{E_{cr}}$, such that a black hole with $M>M_{cr}$ will discharge very slowly (as $E_{eh}<E_{cr}$). In this range the black hole would lose its mass via Hawking radiation until it reaches extremality. On the other hand, a black hole with $M<M_{cr}$ will discharge while approximately keeping $E_{eh}=E_{cr}$, that is, $\frac{Q}{M}\sim ME_{cr}$. This means that in this range, $\frac{Q}{M}$ decreases roughly linearly with $M$. This varying $\frac{Q}{M}$ would result in a more complicated dynamics inside the evaporating black hole, which we will here avoid for the sake of simplicity.
One of the main reasons we are interested in charged black holes is as a simplified toy model for spinning black holes, because the charge $Q$ serves a similar role as the spin parameter $a$ with regards to the formation of a Cauchy horizon inside the black hole. In the case of a spinning black hole, Page [27] found that the angular momentum of the black hole would evaporate at a significantly faster rate than its mass. In other words, for evaporating spinning black hole, $\frac{a}{M}$ decreases to zero during the evaporation process. Thus, from this perspective too, our assumption of $\frac{Q}{M}=const$ is again not realistic. Despite this assumption being unrealistic, we still choose it here in order to simplify the evaporation model and its analysis.
We also need to choose the function $\delta Q$. We’ll choose it in the following form:
$\delta Q=\log\left(1+e^{\frac{u+v}{M(v)}}\right)M(v)^{-1}\varepsilon_{0}M_{0}^{2}\alpha$ (3.22)
This structure of $\delta Q$ contributes currents flowing in the $r_{*}=\frac{u+v}{2}$ direction (at leading order in $\varepsilon_{0}$). $\alpha$ is a free dimensionless parameter controlling the strength of the semiclassical currents, which we typically choose to be of order unity. For the dependence of $\delta Q$ on $u+v$, we chose it to be via the function $f(x)=\log\left(1+e^{x}\right)$ as it interpolates smoothly between $f(x)=0$ at $x\ll 0$ (as desired at the EH) and $f(x)=x$ at $x\gg 0$ (corresponding to the region near the IH). The argument of this function $f(x)$ is $x=\frac{u+v}{M(v)}$, which is a dimensionless variable proportional to $r_{*}=\frac{u+v}{2}$. Since $\delta Q$ is initiated from semiclassical currents, it should be proportional to $\hbar$ ($\propto\varepsilon_{0}M_{0}^{2}$). The later has mass squared dimensions, and the dimensions of $Q$ is $M^{1}$, therefore we also need to multiply the expression by $M(v)^{-1}$ to get the correct mass dimensions.
Note that we use $M(v)$ to get the correct mass dimensions, and not $M_{0}$. Because the drift in the mass is extremely slow, at every point we have a very good local RN approximation. We therefore may naturally expect that the semiclassical currents – and therefore also $\delta Q$ – will be well approximated by the corresponding semiclassical quantities on RN background. From dimensional arguments, semiclassical $\delta Q$ on RN background should take the form $\hbar M^{-1}H_{\delta Q}(\frac{r}{M},\frac{Q}{M})$. We assume that this form also holds in the evaporation case, as explained above, and we also substitute $\hbar=\left[3H_{M}\left(\frac{Q}{M}\right)^{-1}\right]\varepsilon_{0}M_{0}^{2}$. For reasons explained above, we chose $H_{\delta Q}$ to be of the form corresponding to Eq. (3.22). It should be emphasized that $M_{0}$ only appears (squared) along with $\varepsilon_{0}$ as a replacement for $\hbar$.
The main motivation for introducing $\delta Q$ is to demonstrate numerically that semiclassical drift in the charge $Q$ that may take place and even become significant near the IH does not affect the backreaction of the metric near the IH – once the four input functions $M(v)$, $Q^{v}(v)$, $\widetilde{T}_{uu}^{(-)}(v)$, $\text{$\widetilde{T}_{vv}^{(-)}$}(v)$ have been specified. (Note that $\delta Q$ may have an effect on the value of $\widetilde{T}_{uu}^{(-)}(v)$, $\text{$\widetilde{T}_{vv}^{(-)}$}(v)$, similar to that of the sources $Z_{R}$$Z_{S}$.) This fact is already embodied in the analytical approximation – where the charge function required is $Q^{v}(v)$ only. Note that in the absence of $\delta Q$, from Eq. (2.27) $J_{u}=0$. Non-vanishing $\delta Q$ allows us to explore scenarios with $J_{u}$ as well. Note that $\delta Q$ is just the simplest model we could choose which will model current $J_{u}$.
3.6 Choice of source functions $Z_{R}$, $Z_{S}$
As already mentioned in subsection 2.8, for the task of validating the analytical approximation, we can freely choose the semiclassical functions which are $T_{uv}^{(SC)}$ and $T_{\theta\theta}^{(SC)}$. This amounts to freely choosing the two source functions $Z_{R}$, $Z_{S}$ defined in Eqs. (2.1172.118).
We find it convenient to choose these source functions in the following form:
$Z_{R}=\varepsilon_{0}M_{0}^{2}M^{-3}e^{S}L_{0}\left(Z_{R0}+\frac{r}{M}Z_{Rr}\right)$ (3.23)
$Z_{S}=\varepsilon_{0}M_{0}^{2}M^{-5}e^{S}L_{0}\left(Z_{S0}+\frac{r}{M}Z_{Sr}\right)$ (3.24)
There are 4 parameters: $Z_{R0}$$Z_{Rr}$$Z_{S0}$$Z_{Sr}$, all dimensionless.
The reasoning behind this choice is as follows. In the semiclassical theory these source functions vanish both at the EH and the IH. We therefore choose these functions to be proportional to $e^{S}$, which is close to zero near both the EH and IH [^5] For RN background in Unruh state, it was found [30] that near the IH $Z_{R}$ is proportional to $e^{S}\sim e^{-2\kappa_{-}r_{*}}$, whereas $Z_{S}$ decays as $r_{*}^{-5}$. In our numerical simulation, we decided to simplify the computation scheme by assuming both to be proportional to $e^{S}$ near the IH. We do not expect this to change the resultant backreaction metric because $r_{*}^{-5}$ vanishes sufficiently fast.. We then multiply $e^{S}$ by the length scale $L_{0}$ similar to the way it appears in the evolution equations, see Eqs. (2.37,2.38). The source term is semiclassical, therefore it should be proportional to $\hbar$ ($\propto\varepsilon_{0}M_{0}^{2}$) which has mass squared dimensions, so we want to create a dimensionless quantity from that, leading us to the dimensionless factor $\hbar M^{-2}\propto\varepsilon_{0}M_{0}^{2}M^{-2}$. Since $F_{1}(R)]=\left[\frac{1}{\sqrt{R}}\right]=[l]^{-1}$, the $Z_{R}$ source term is overall multiplied by $M^{-3}$, and since $F_{2}(R)]=\left[\frac{1}{R^{\frac{3}{2}}}\right]=[l]^{-3}$, the $Z_{S}$ source term is overall multiplied by $M^{-5}$. For convenience, we chose the spacetime dependence of these sources to be dependent on $\frac{r}{M}$ only, and we further restricted our choice to linear functions of this variable, determined by the four arbitrary parameters $Z_{R0}$$Z_{Rr}$$Z_{S0}$$Z_{Sr}$. When running our numerical code, we’ll choose multiple specific (arbitrary) combinations of these parameters and the charge parameter $\alpha$.
The presence of these source terms, and / or the changing charge, cause non-conservation of $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$, as already mentioned in subsection 2.8. We shall now calculate the amount of this non-conservation, using the evolution equations and the source terms in Eqs. (3.23,3.24). We start by taking the derivative of the constraint Eq. (2.43):
$\left(\widetilde{T}_{uu}\right)_{,v}$ $=R_{,uv}S_{,u}+R_{,u}S_{,uv}-\left(R_{,uv}\right)_{,u}$ (3.25)
$=L_{0}e^{S}\left(F_{1}(R,Q)+\varepsilon_{0}M^{-3}\left(Z_{R0}+\frac{r}{M}Z_{Rr}\right)\right)S_{,u}$ (3.26)
$+\left(e^{S}L_{0}F_{2}(R,Q)+\varepsilon_{0}M^{-5}e^{S}L_{0}\left(Z_{S0}+\frac{r}{M}Z_{Sr}\right)\right)R_{,u}$ (3.27)
$-\left[L_{0}e^{S}\left(F_{1}(R,Q)+\varepsilon_{0}M^{-3}\left(Z_{R0}+\frac{r}{M}Z_{Rr}\right)\right)\right]_{,u}$ (3.28)
$=L_{0}e^{S}\left(F_{1}(R,Q)+\varepsilon_{0}M^{-3}\left(Z_{R0}+\frac{r}{M}Z_{Rr}\right)\right)S_{,u}$ (3.29)
$+\left(e^{S}L_{0}F_{2}(R,Q)+\varepsilon_{0}M^{-5}e^{S}L_{0}\left(Z_{S0}+\frac{r}{M}Z_{Sr}\right)\right)R_{,u}$ (3.30)
$-L_{0}e^{S}\left(F_{1}(R,Q)+\varepsilon_{0}M^{-3}\left(Z_{R0}+\frac{r}{M}Z_{Rr}\right)\right)S_{,u}$ (3.31)
$-e^{S}L_{0}\frac{\partial F_{1}(R,Q)}{\partial R}R_{,u}-e^{S}L_{0}\frac{\partial F_{1}(R,Q)}{\partial Q}Q_{,u}-\varepsilon_{0}M^{-3}e^{S}L_{0}\frac{r_{,u}}{M}Z_{Rr}$ (3.32)
$=e^{S}L_{0}\left[\left(\varepsilon_{0}M^{-5}\left(Z_{S0}+\frac{r}{M}Z_{Sr}\right)\right)R_{,u}-\frac{2Q}{R^{\frac{3}{2}}}Q_{,u}-\varepsilon_{0}M^{-3}\frac{r_{,u}}{M}Z_{Rr}\right]$ (3.33)
And similarly for $\widetilde{T}_{vv}:$
$\left(\widetilde{T}_{vv}\right)_{,u}=e^{S}L_{0}\left[\left(\varepsilon_{0}M^{-5}\left(Z_{S0}+\frac{r}{M}Z_{Sr}\right)\right)R_{,v}-\frac{2Q}{R^{\frac{3}{2}}}Q_{,v}-\varepsilon_{0}M^{-3}\frac{r_{,v}}{M}Z_{Rr}\right]$ (3.34)
From Eq. (3.33) we can see that if all sources are zero and in addition $Q_{,u}=0$, then $\widetilde{T}_{uu,v}=0$. Since $T_{uu}$ vanishes on the initial hypersurface, it follows that in such a case, $T_{uu}$ vanishes in the entire domain of dependence. (Also notice that the parameter $Z_{R0}$ does not affect the conservation of either $\widetilde{T}_{uu}$ or $\widetilde{T}_{vv}$.)
As was mentioned in the Introduction, the analytical approximation is based on a key assumption about the behavior of the RSET in the evolving geometry. Here we shall specify this assumption. In the near-IH RN background (in Unruh state) it was found that both $Z_{R}$ and $Z_{S}$ [30] vanish sufficiently fast in both $u$ and $v$ and hence become negligibly small on approaching the IH (see also footnote 5). Our key assumption is that in the evolving geometry too, these source functions remain negligibly small. This assumption leads to conservation of the fluxes $\widetilde{T}_{uu}$ and $\widetilde{T}_{vv}$ in the near-IH region: $\widetilde{T}_{uu,v}=0$ and $\widetilde{T}_{vv,u}=0$. To see this, we evaluate $\widetilde{T}_{uu,v}$ by taking the $v$ derivative of Eq. (2.43) and also substituting the evolution equations Eqs. (2.115,2.116):
$\left(\widetilde{T}_{uu}\right)_{,v}$ $=R_{,uv}S_{,u}+R_{,u}S_{,uv}-\left(R_{,uv}\right)_{,u}$ (3.35)
$=\left(e^{S}L_{0}F_{1}(R)+Z_{R}\right)S_{,u}$ (3.36)
$+\left(e^{S}L_{0}F_{2}(R)+Z_{S}\right)R_{,u}$ (3.37)
$-\left(e^{S}L_{0}F_{1}(R)+Z_{R}\right)_{,u}$ (3.38)
Since $e^{S}$ decays exponentially with $u$ and $v$ (see subsection 4.5), the quick vanishing of $Z_{R}$ and $Z_{S}$ leads to $\left(\widetilde{T}_{uu}\right)_{,v}=0$ in the near-IH region. The same argument also applies to $\left(\widetilde{T}_{vv}\right)_{,u}$.
4 Analytical approximation for the near-IH metric inside slowly evaporating black holes
In slowly evaporating black holes, the evaporation rate
$\varepsilon=M_{,v}$ (4.1)
is small, $\varepsilon\ll 1$. We can build an approximation assuming $\varepsilon$ is small. Note that this $\varepsilon$ is distinct from $\varepsilon_{0}$: $\varepsilon$ depends on $v$ while $\varepsilon_{0}$ is a constant associated with the substitution $v=v_{0}$. This difference reflects the fact that at very late $v$, $\varepsilon$ is no longer small, being proportional to $M(v)^{-2}$.
4.1 Regions inside black hole
Figure 4.1: metrics inside black hole
We divide the space-time domain inside the black hole to three regions, as seen in Figure 4.1 . Those regions, numbered 1-3, are ordered in chronological order. Each two successive regions overlap in their area of validity, and together they cover the entire interior of the black hole up to the inner horizon.
Region 1 is the neighborhood of the event horizon. This is taken here to be ingoing charged Vaidya metric Eq. (2.99), which we later convert into double-null coordinates.
Region 2 is between the event horizon and the inner horizon (but not too close to the latter). This region is characterized by local RN approximation: at each point in that region, the metric is well approximated by the RN metric (up to deviations of order $\varepsilon$), though with slowly varying parameters of mass $M(v)$ and charge $Q(v)$.
It is important to note that in region 2, the local RN parameters $M$ and $Q$ are constant along ingoing constant-$v$ rays, as these parameters depend only on $v$ but not on $u$. Also note that region 2 completely covers region 1.
Region 3 is the region to the future of region 2 (although there is some overlap between these two regions) where $r_{*}\gg M$. This means that, within the local RN approximation, we are placed very close to the IH. Note, however, that for $u+v$ which are extremely large – comparable to the black hole evaporation time – the local RN approximation no longer hold (due to drift in $R$, see below). For this reason, region 3 is not entirely contained in region 2.
As mentioned above, in region 2 the local RN parameter $Q$ depends on $v$ but not on $u$. This happens because of the following reason: at the event horizon $J_{u}=0$ as dictated by regularity reasons, while $J_{v}$ need not vanish at the horizon. This regularity reason is as follows: the regular Kruskal coordinate at the event horizon is $U=e^{\kappa_{+}u}$, therefore $u=\frac{\ln U}{\kappa_{+}}$. The quantity $J_{U}=\frac{\partial u}{\partial U}J_{u}=\frac{1}{U\kappa_{+}}J_{u}$ must be finite, however the event horizon is at $U=0$, which makes this quantity diverge unless $J_{u}=0$. So this regularity reason dictate that $J_{u}=0$. For similar regularity reasons, $T_{uu}=0$ at the EH.
4.2 Basic ansatz of the analytical approximation for region 3
Consider first the near-IH domain in region 2 where the geometry is locally RN. We saw in Eq. (2.89) that in RN $S_{,u}=S_{,v}=-\kappa_{-}$ near the IH. We therefore get that $S$ is monotonically decreasing in region 2, and becomes very negative upon increasing $u$ and $v$. Therefore, $e^{S}$ becomes negligibly small at large $u+v$.
For the analysis of region 3, we start by assuming that the behavior of $S$ in region 3 is similar to region 2 near the IH in the following sense: $S$ continues to decrease monotonically with increasing $u$ and with $v$; as a consequence, $e^{S}$ that was already negligible in region 2 near IH, will remain negligible throughout region 3. We justify this ansatz later.
Using the evolution equations Eq. (2.115,2.116) and neglecting the factor $e^{S}$ – and also assuming negligibly small sources $Z_{R}\approx Z_{S}\approx 0$ (see subsection 3.6) – we get, throughout region 3:
$R_{,uv}=0,\quad S_{,uv}=0$ (4.2)
Therefore both functions $R$,$S$ in region 3 are functions of $v$ plus functions of $u$:
$R(u,v)=R^{v}(v)+R^{u}(u)$ (4.3)
$S(u,v)=S^{v}(v)+S^{u}(u)$ (4.4)
Because of that, $S_{,v}$ and $R_{,v}$ are functions of $v$ only: $S_{,v}(v),R_{,v}(v).$ Similarly, $S_{,u}$ and $R_{,u}$ are functions of $u$ only: $S_{,u}(u),R_{,u}(u)$. We still need to determine those four functions, which is our next task.
Figure 4.2: Diagram explaining analytical approximation
Region 3 is between CH and the light-blue line numbered 3. Region 2 is between EH and the red line numbered 2.
4.3 Computing $S_{,v}$
Refer to Figure 4.2. Consider a point $P$ deep in region 3. We first focus on computing $S_{,v}$ at point $P$. Since $S_{,v}(v)$ is constant along the red constant-$v$ ray, we can read its value from point $P_{v}$, which is in the overlap of regions 2 and 3, and therefore its value is obtained from the corresponding expression in region 2 near IH (see Eq. (2.89)):
$S_{,v}=-\kappa_{-}^{v}(v)$ (4.5)
where $\kappa_{-}^{v}(v)$ refers to the value of $\kappa_{-}$ in the local RN at point $P_{v}$:
$\kappa_{-}^{v}(v)\equiv\kappa_{-}(M(v),Q^{v}(v))$ (4.6)
4.4 Computing $S_{,u}$
We next move to analyze $S_{,u}(u)$ at point $P$ in region 3. This quantity is constant along the constant-$u$ blue ray, so that its value may be obtained by carrying this blue ray until it enters the overlap between region 3 and near-IH region 2, say until the point marked $P_{u}$:
$S_{,u}(P)=-\kappa_{-}(P_{u})$ (4.7)
Here $\kappa_{-}(P_{u})$ denotes the value of $\kappa_{-}$ at the local RN at point $P_{u}$. We pick the point $P_{u}$ in the matching zone between region 2 and region 3. In this matching zone, both approximation of region 2 and region 3 should apply. The approximation of region 3 requires $e^{S}$ to be negligible. The region 2 approximation requires that the near IH drift in $R$ is still negligible, that is $\varepsilon\cdot\left(v+u\right)\ll M$. We choose this matching zone to be at the region where $u+v$ is of order, say, $\frac{1}{\kappa_{-}}\log\varepsilon$. At that point, $e^{S}$ is indeed negligible, because $S\approx-\kappa_{-}(u+v)$ so $e^{S}\approx\varepsilon\ll 1$. In addition, $\varepsilon\cdot\left(v+u\right)=-\frac{1}{\kappa_{-}}\varepsilon\log\varepsilon$ is negligible compared to $M$ in that zone.
As mentioned above, the local RN at a given point in region 2 is the same along the constant-$v$ ray that passes through that point. We could therefore mark a constant-$v$ ray from this point $P_{u}$ up to the point $P_{2}$ on the EH. However we prefer to pick another point $P_{1}$ on the EH to represent the local RN at point $P_{u}$. This point $P_{1}$ is obtained by continuing from point $P_{u}$ along the dashed blue constant-$u$ ray until we meet the hypersurface $r=M$ (which is also the $r_{*}=0$ line) at point $P_{0}$. Then we move from $P_{0}$ along the dashed purple constant-$v$ ray to reach the point $P_{1}$ on the EH. We prefer to work with the representative point $P_{1}$ rather than $P_{2}$ because this choice leads to a significantly simpler expression for $S_{,u}$ (and $R_{,u}$) in region 3. However, it remains to be explained why this choice, namely the replacement of point $P_{2}$ by point $P_{1}$, is justified.
[…]
Since at point $P_{0}$ the coordinates $u$,$v$ satisfy $r_{*}=\frac{u+v}{2}=0$, we get $u_{P}=u_{P_{0}}=-v_{P_{0}}=-v_{P_{1}}$(where $u_{P}$ is $u$ at point $P$, etc.). As explained above, in order to compute the value of $\kappa_{-}$ at point $P_{u}$, we take the values of $M$,$Q$ according to the $v$ coordinate of point $P_{1}$ (which is equal to that of $P_{0}$). We are therefore interested in the quantity:
$M(v_{P_{1}})=M(v=-u_{P_{0}})=M(v=-u_{P})$ (4.8)
This motivates us to define the function $M^{u}(u)$ as follows:
$M^{u}(u)\equiv M(v=-u)$ (4.9)
Similarly for $Q$:
$Q^{u}(u)\equiv Q^{v}(v=-u)$ (4.10)
According to the same lines, we also define
$\kappa_{-}^{u}(u)\equiv\kappa_{-}(M^{u}(u),Q^{u}(u))$ (4.11)
or in other words,
$\kappa_{-}^{u}(u)\equiv\kappa_{-}(v=-u)$ (4.12)
For later use, we similarly define the $\kappa_{+}$ version of this quantity:
$\kappa_{+}^{u}(u)\equiv\kappa_{+}(v=-u)$ (4.13)
Summarizing these notational relations, we get (recalling that $u_{P}=u_{P_{u}}$):
$\kappa_{-}(P_{u})=\kappa_{-}^{u}(u_{P_{u}})=\kappa_{-}^{u}(u_{P})$ (4.14)
We can therefore rewrite Eq. (4.7) as
$S_{,u}(P)=-\kappa_{-}^{u}(u_{P})$ (4.15)
Or in other words, for any point in region 3:
$S_{,u}=-\kappa_{-}^{u}(u)$ (4.16)
4.5 Justifying the basic ansatz
Recalling Eqs. (4.5,4.16) and the fact that $\kappa_{-}>0$, we see that in region 3, $S$ continues to decrease monotonically with both $u$ and $v$. Since $e^{S}$ was already exponentially small in the overlap between region 2 and region 3, it is totally negligible throughout region 3 – confirming our former ansatz about this issue.
Correspondingly we can neglect the factor $e^{S}$ also in (3.33,3.34) and we get that:
$\widetilde{T}_{uu,v}=\widetilde{T}_{vv,u}=0$ (4.17)
Therefore in region 3 $\widetilde{T}_{vv}$ is a function of $v$ and $\widetilde{T}_{uu}$ is a function of $u$.
4.6 Computing $R_{,v}$
We next focus on $R_{,v}$ in region 3. We recall that $\widetilde{T}_{vv}\equiv 8\pi r^{2}T_{vv}=R_{,v}S_{,v}-R_{,vv}$, and we make an ansatz that $R_{,vv}\ll R_{,v}S_{,v}$ (to be justified later) and therefore, using Eq. (4.5) and neglecting $R_{,vv}$, we get:
$R_{,v}=\frac{\widetilde{T}_{vv}}{S_{,v}}=-\frac{\widetilde{T}_{vv}^{(-)}}{\kappa_{-}(v)}$ (4.18)
We assume that $\widetilde{T}_{vv}\propto\varepsilon$, therefore $R_{,v}\propto\varepsilon$.
We proceed now to show that $R_{,vv}\propto\varepsilon^{2}$ and therefore it is indeed negligible relative to $R_{,v}S_{,v}$. We start with a simple intuitive argument: $R_{,v}=-\frac{\widetilde{T}_{vv}^{(-)}}{\kappa_{-}(v)}$ is already of order $\varepsilon$. Therefore $R_{,vv}$ would exclusively be made of contributions from the $v$ derivative of either $\kappa_{-}(v)$ or $\widetilde{T}_{vv}^{(-)}$ . But since both these quantities are uniquely determined by $M$ and $Q^{v}$, their $v$ derivative only originates from the $v$ derivatives of $M$ and $Q^{v}$. These $v$ derivatives are obviously proportional to $\varepsilon$, and therefore $R_{,vv}\propto\varepsilon^{2}$.
We shall also give here a more systematic derivation of this property. Let $F$ be a quantity that can be expressed as:
$F=M^{m}\varepsilon^{n}H_{F}\left(\frac{Q^{v}(v)}{M(v)}\right)$ (4.19)
for some dimensionless function $H_{F}$ that doesn’t depend on $\hbar$. It’s important to note that $H_{F}$ doesn’t depend on the position in any way other than via the dependence of $Q^{v}$ and $M$ on $v$. (All the functions $H_{i}\left(\frac{Q}{M}\right)$ that will appear below will all be of this class: dimensionless and won’t depend on $\hbar$, and won’t depend on the position except via $Q^{v}(v)$,$M(v)$) Then we’ll prove that:
$F_{,v}=M^{m-1}\varepsilon^{n+1}H_{F^{\prime}}\left(\frac{Q^{v}(v)}{M(v)}\right)$ (4.20)
for some new dimensionless function $H_{F^{\prime}}$. For that purpose, we calculate $\left(\frac{Q^{v}}{M}\right)_{,v}:$
$\left(\frac{Q^{v}}{M}\right)_{,v}=-\frac{Q^{v}M_{,v}}{M^{2}}+\frac{Q_{,v}^{v}}{M}$ (4.21)
From semiclassical theory we know that Eq. (3.7) (setting $M^{(+)}\rightarrow M$ and $Q\rightarrow Q^{v}$):
$M_{,v}\equiv-\varepsilon=-\frac{\hbar}{M^{2}}H_{M}\left(\frac{Q^{v}}{M}\right)$ (4.22)
and similarly for $Q_{,v}^{v}$:
$Q_{,v}^{v}=-\frac{\hbar}{M^{2}}H_{Q}\left(\frac{Q^{v}}{M}\right)=-\varepsilon\frac{H_{Q}\left(\frac{Q^{v}}{M}\right)}{H_{M}\left(\frac{Q^{v}}{M}\right)}\equiv-\varepsilon H_{q}\left(\frac{Q^{v}}{M}\right)$ (4.23)
Therefore,
$\left(\frac{Q^{v}}{M}\right)_{,v}=\frac{\varepsilon Q^{v}}{M^{2}}-\frac{\varepsilon H_{q}\left(\frac{Q^{v}}{M}\right)}{M}\equiv\frac{\varepsilon}{M}H_{v}\left(\frac{Q^{v}}{M}\right)$ (4.24)
We can now use this result to calculate $\frac{d}{dv}H\left(\frac{Q^{v}}{M}\right)$:
$\frac{d}{dv}H\left(\frac{Q^{v}}{M}\right)=H^{\prime}\left(\frac{Q^{v}}{M}\right)\frac{d}{dv}\left(\frac{Q^{v}}{M}\right)=H^{\prime}\left(\frac{Q^{v}}{M}\right)\frac{\varepsilon}{M}H_{v}\left(\frac{Q^{v}}{M}\right)$ (4.25)
where $H^{\prime}$ denotes the derivative of the function $H$ with respect to its argument $\frac{Q^{v}}{M}$. We next calculate $\varepsilon_{,v}$ by differentiating Eq. (4.22):
$\varepsilon_{,v}$ $=-2\frac{\hbar M_{,v}}{M^{3}}H_{M}\left(\frac{Q^{v}}{M}\right)+\frac{\hbar}{M^{2}}\frac{d}{dv}H_{M}\left(\frac{Q^{v}}{M}\right)$ (4.26)
$=-2\frac{\hbar M_{,v}}{M^{3}}H_{M}\left(\frac{Q^{v}}{M}\right)+\frac{\hbar}{M^{2}}H^{\prime}_{M}\left(\frac{Q^{v}}{M}\right)\frac{\varepsilon}{M}H_{v}\left(\frac{Q^{v}}{M}\right)$ (4.27)
$=\frac{\hbar}{M^{2}}\frac{\varepsilon}{M}\left[2H_{M}\left(\frac{Q^{v}}{M}\right)+H^{\prime}_{M}\left(\frac{Q^{v}}{M}\right)H_{v}\left(\frac{Q^{v}}{M}\right)\right]$ (4.28)
$\equiv\frac{\varepsilon^{2}}{M}H_{\varepsilon}\left(\frac{Q^{v}}{M}\right)$ (4.29)
where $H^{\prime}_{M}$ denotes the derivative of the function $H_{M}$ with respect to its argument $\frac{Q^{v}}{M}$. Finally, we take the derivative of $F$ in Eq. (4.19) to get:
$F_{,v}$ $=mM_{,v}M^{m-1}\varepsilon^{n}H_{F}\left(\frac{Q^{v}}{M}\right)+n\varepsilon_{,v}M^{m}\varepsilon^{n-1}H_{F}\left(\frac{Q^{v}}{M}\right)+M^{m}\varepsilon^{n}H^{\prime}_{F}\left(\frac{Q^{v}}{M}\right)\frac{\varepsilon}{M}H_{v}\left(\frac{Q^{v}}{M}\right)$ (4.30)
$=-M^{m-1}\varepsilon^{n+1}mH_{F}\left(\frac{Q^{v}}{M}\right)+M^{m-1}\varepsilon^{n+1}nH_{\varepsilon}\left(\frac{Q^{v}}{M}\right)H_{F}\left(\frac{Q^{v}}{M}\right)$ (4.31)
$+M^{m-1}\varepsilon^{n+1}H^{\prime}_{F}\left(\frac{Q^{v}}{M}\right)H_{v}\left(\frac{Q^{v}}{M}\right)$ (4.32)
$\equiv M^{m-1}\varepsilon^{n+1}H_{F^{\prime}}\left(\frac{Q^{v}}{M}\right)$ (4.33)
We proved Eq. (4.20). Now we will use this result for Eq. (4.18). First, recall from Eq. (3.17) that $\widetilde{T}_{vv}^{(-)}$ can be expressed as:
$\widetilde{T}_{vv}^{(-)}=\varepsilon H_{T_{vv}}^{(-)}\left(\frac{Q^{v}}{M}\right)$ (4.34)
Secondly, recall that for $\kappa_{-},$Eq. (2.53) can be written as:
$\kappa_{-}=\frac{1}{M}H_{\kappa}\left(\frac{Q^{v}}{M}\right)$ (4.35)
Therefore,
$R_{,v}=-\frac{\widetilde{T}_{vv}^{(-)}}{\kappa_{-}(v)}=-\varepsilon M\frac{H_{T_{vv}}^{(-)}\left(\frac{Q^{v}}{M}\right)}{H_{\kappa}\left(\frac{Q^{v}}{M}\right)}\equiv\varepsilon MH_{R}\left(\frac{Q^{v}}{M}\right)$ (4.36)
We can now apply Eq. (4.20) to $F=R_{,v}$ to get $R_{,vv}$:
$R_{,vv}=\varepsilon^{2}H_{vv}\left(\frac{Q^{v}}{M}\right)$ (4.37)
We also express $R_{,v}S_{,v}$ in a similar format:
$R_{,v}S_{,v}$ $=-R_{,v}\kappa_{-}=\left[\varepsilon MH_{R}\left(\frac{Q^{v}}{M}\right)\right]\left[\frac{1}{M}H_{\kappa}\left(\frac{Q^{v}}{M}\right)\right]$ (4.38)
$=\varepsilon H_{RS}\left(\frac{Q^{v}}{M}\right)$ (4.39)
This confirms our ansatz that $R_{,vv}\ll R_{,v}S_{,v}$, as $R_{,v}\propto\varepsilon$, while $R_{,vv}\propto\varepsilon^{2}$. Our final result is therefore, throughout region 3,
$R_{,v}(u,v)=-\frac{\widetilde{T}_{vv}^{(-)}(v)}{\kappa_{-}^{v}(v)}$ (4.40)
4.7 Computing $R_{,u}$
We next focus on $R_{,u}$. We recall that $\widetilde{T}_{uu}\equiv 8\pi r^{2}T_{uu}=R_{,u}S_{,u}-R_{,uu}$, and we make a similar ansatz that $R_{,uu}\ll R_{,u}S_{,u}$ because $R_{,uu}\propto\varepsilon^{2}$ while $R_{,u}\propto\varepsilon$ (justifying it is similar to the $R_{,v}$ case discussed above). Therefore, using Eq. (4.16), and recalling that in region 3, $\widetilde{T}_{uu}=\widetilde{T}_{uu}^{(-)}(v=-u)$, we get:
$R_{,u}(u,v)=-\frac{\widetilde{T}_{uu}^{(-)}(v=-u)}{\kappa_{-}^{u}(u)}$ (4.41)
To avoid confusion we should note the following: $\widetilde{T}_{uu}^{(-)}$ was defined as a function of $v$ (as the near-IH $\widetilde{T}_{uu}$ value in RN with parameters $M(v)$ and $Q(v)$). We also saw that in region 3, $\widetilde{T}_{uu}$ depends only on $u$ (see subsection 4.5). The expression we used above $\widetilde{T}_{uu}=\widetilde{T}_{uu}^{(-)}(v=-u)$ is indeed consistent with both of these requirements.
4.8 Summary
Eqs. (4.5,4.16,4.40,4.41) describe the analytical approximation for the quantities $S_{,v}$, $S_{,u}$, $R_{,v}$, $R_{,u}$ in region 3. As can be seen in these equations, this analytical approximation only depends on a set of four semiclassical input functions: $M(v)$ and $Q^{v}(v)$ (or alternatively $\widetilde{T}_{vv}^{(+)}(v)$ and $J_{v}^{(+)}(v)$) at the event horizon, and $\widetilde{T}_{uu}^{(-)}(v)$, $\text{$\widetilde{T}_{vv}^{(-)}$}(v)$ at the inner horizon. (Recall, these two last quantities are defined, via the semiclassical theory, as the near-IH fluxes in RN with parameters $M(v)$ and $Q(v)$, and are therefore viewed as functions of $v$.)
5 Numerical algorithm
We would like to integrate the evolution equations Eq. (2.37,2.38) numerically. We do that by putting both $R$ and $S$ on a discrete grid. We denote the distance between grid points in the $u$ direction as $du$ and the distance between grid points in the $v$ direction as $dv$. The grid points are accessed by two integer indices $i,j$. For simplicity we chose $du=dv$ throughout this work.
The basic numerical algorithm computes $R$ at the point $(i,j)$ (marked red in Figure 5.1) as a combination of 3 previous points (marked blue), and of the derivative $R_{,uv}$ evaluated at the middle point $(i-\nicefrac{{1}}{{2}},j-\nicefrac{{1}}{{2}})$ (marked green), as illustrated in Figure 5.1, as follows:
Figure 5.1: Points in numerical algorithm
$R(i,j)=R(i-1,j)+R(i,j-1)-R(i-1,j-1)+du\;dvR_{,uv}(i-\nicefrac{{1}}{{2}},j-\nicefrac{{1}}{{2}})$ (5.1)
A similar calculation is done for $S$:
$S(i,j)=S(i-1,j)+S(i,j-1)-S(i-1,j-1)+du\;dvS_{,uv}(i-\nicefrac{{1}}{{2}},j-\nicefrac{{1}}{{2}})$ (5.2)
Note that the green point $(i-\nicefrac{{1}}{{2}},j-\nicefrac{{1}}{{2}})$ is not a grid point, as indicated by the half integer indices. To evaluate quantities at that point, we take the average of the relevant quantities in the two neighboring points (i1,j)(i-1,j) and (i,j1)(i,j-1).
Such a discrete numerical algorithm necessarily introduces a numerical error, called truncation error, which in this specific algorithm should be proportional to $du\cdot dv=du^{2}$. We can measure this numerical error by running the algorithm with two different values of $du$ and comparing the results obtained in these two runs, see subsection 9.1.
6 Initial values
In the vicinity of the EH, we presumably have, to a good approximation, a charged Vaidya solution, which evaporates at the rate of $\varepsilon$. Recall Eq. (2.99):
$ds^{2}=-\left(1-\frac{2M(v)}{r}+\frac{Q(v)^{2}}{r^{2}}\right)dv^{2}+2dvdr+r^{2}d\Omega^{2}$ (6.1)
In order to connect to the formalism developed above in subsection 2.1 we have to transform it into double null coordinates. Unfortunately, we cannot implement this transformation analytically. Nevertheless, since $\varepsilon\ll 1$, at each point in this Vaidya geometry there is a local approximately-RN solution. Our strategy will be to use this fact to construct the (approximate) metric for the Vaidya geometry in double null coordinates.
6.1 “naive $g_{1}$ metric”
We choose a certain $v$ value which we denote $v_{0}$. Due to the smallness of $\varepsilon$ there is a range of $v$ in the vicinity of $v_{0}$ in which the geometry is RN to a good approximation. The Vaidya metric there reduces to the RN geometry in ingoing Eddington coordinates $v,r$. We now transform it to double null coordinates $u,v$. Taking Eq. (2.79) and applying with local $M(v_{0})$ and $Q(v_{0})$ parameters:
$ds^{2}=\left(1-\frac{2M(v_{0})}{r(u,v)}+\frac{Q(v_{0})^{2}}{r(u,v)^{2}}\right)du\;dv+r(u,v)^{2}d\Omega^{2}$ (6.2)
where $r(u,v)$ is a function of $\frac{u+v}{2}=r_{*}$ given implicitly by the relation:
$r_{*}=r-M(v_{0})+\frac{1}{2\kappa_{+}(v_{0})}\log\left|\frac{r-r_{+}(v_{0})}{M(v_{0})-r_{+}(v_{0})}\right|-\frac{1}{2\kappa_{-}(v_{0})}\log\left|\frac{r-r_{-}(v_{0})}{M(v_{0})-r_{-}(v_{0})}\right|$ (6.3)
We will now seek to construct a global metric that applies in the entire range of $v$. To achieve this we will simply replace the parameter $v_{0}$ with $v$, resulting in the following metric which we denote “naive $g_{1}$ metric” (for reasons that will be clarified soon):
$ds^{2}=\left(1-\frac{2M(v)}{r(u,v)}+\frac{Q(v)^{2}}{r(u,v)^{2}}\right)du\;dv+r(u,v)^{2}d\Omega^{2}$ (6.4)
The equation defining $r(u,v)$ then becomes:
$\frac{u+v}{2}\equiv r_{*}=r-M(v)+\frac{1}{2\kappa_{+}(v)}\log\left|\frac{r-r_{+}(v)}{M(v)-r_{+}(v)}\right|-\frac{1}{2\kappa_{-}(v)}\log\left|\frac{r-r_{-}(v)}{M(v)-r_{-}(v)}\right|$ (6.5)
For later use, we define $f$ which is a function of $v$,$r$:
$f=1-\frac{2M(v)}{r}+\frac{Q(v)^{2}}{r^{2}}$ (6.6)
Note that
$\left(\frac{\partial r}{\partial r_{*}}\right)_{v}=f$ (6.7)
The metric in Eq. (6.4) provides a description of the Vaidya metric Eq. (6.1) in double null coordinates, up to deviation of order $\varepsilon\ll 1$. We shall now discuss the regularity of this metric at the EH.
Already in RN, the metric in $u$,$v$ Eddington coordinates isn’t regular at the EH, as $g_{uv}=0$ and hence the metric has zero determinant there. Nevertheless, in RN we can define double null Kruskal coordinate $U(u)$ and $V(v)$ with the transformation
$U=e^{\kappa_{+}u},\hfill V=e^{\kappa_{+}v}$ (6.8)
Attempting to generalize these Kruskal coordinates to our case, since $\kappa_{+}$ is a function of $v$ (unlike the RN case) $U=e^{\kappa_{+}(v)u}$ is not a function of only $u$, and therefore it is not a null coordinate. Even worse, it turns out that the naive metric (6.4) geometry is not regular at the EH: we will show in Appendix A that the Ricci scalar actually diverges there. To solve this problem we will move to the metric $\widetilde{g_{1}}$ (see next subsection).
Note that this problem arose when we moved from $v_{0}$ in the local RN metric Eq. (6.2) to the global $v$-dependent metric (6.4). We need to implement this transition in a different manner to avoid this pitfall, as we will now describe.
6.2 The $\widetilde{g_{1}}$ metric
First, we will take the local Reissner–Nordström metric (at $v=v_{0}$) in double null coordinates Eq. (6.2), and transform it to different double null coordinates $\widetilde{u}(u),\widetilde{v}(v)$ defined by:
$\widetilde{v}\equiv\mathrel{{\mathop{\int}\limits_{0}^{v}}}\kappa_{+}(v_{0})dv,\qquad\widetilde{u}\equiv\mathrel{{\mathop{\int}\limits_{0}^{u}}}\kappa_{+}(v_{0})du$ (6.9)
Here $\kappa_{+}$is still a constant, but that will change soon once we make the parameters vary with $v$ . Then the new metric is:
$ds^{2}=\frac{1}{\kappa_{+}(v_{0})^{2}}\left(1-\frac{2M(v_{0})}{r}+\frac{Q(v_{0})^{2}}{r^{2}}\right)d\widetilde{u}\;d\widetilde{v}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)$ (6.10)
We define $\widetilde{r_{*}}$ correspondingly as:
$\widetilde{r_{*}}=\frac{\widetilde{v}+\widetilde{u}}{2}$ (6.11)
which satisfies:
$\widetilde{r_{*}}=\kappa_{+}(v_{0})r_{*}$ (6.12)
Now we take this metric and make its parameters $M$,$Q$ (and hence, also $\kappa_{+}$) vary with $\widetilde{v}$. Our metric is then:
$ds^{2}=\frac{1}{\kappa_{+}(\widetilde{v})^{2}}\left(1-\frac{2M(\widetilde{v})}{r}+\frac{Q(\widetilde{v})^{2}}{r^{2}}\right)d\widetilde{u}\;d\widetilde{v}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)$ (6.13)
[…]
$\frac{\widetilde{u}+\widetilde{v}}{2}\equiv\widetilde{r_{*}}=\kappa_{+}(v)\left(r-M(v)+\frac{1}{2\kappa_{+}(v)}\log\left|\frac{r-r_{+}(v)}{M(v)-r_{+}(v)}\right|-\frac{1}{2\kappa_{-}(v)}\log\left|\frac{r-r_{-}(v)}{M(v)-r_{-}(v)}\right|\right)$ (6.14)
Note that when $\widetilde{u}+\widetilde{v}=0$, $r=M(v)$.
We shall call this metric “the $\widetilde{g_{1}}$ metric”.
Next we discuss the regularity of this geometry at the EH. This metric, as is, still isn’t regular, because $g_{\widetilde{u}\widetilde{v}}=0$ at the EH. However, like in RN, we can transform it to Kruskal coordinates (identical to those defined in Eq. (6.8)) which now take the form:
$U=e^{\widetilde{u}},\qquad V=e^{\widetilde{v}}$ (6.15)
Recalling that our metric is just like RN along a line of constant $v$, and recalling Eq. (2.98) , the metric function $g_{\widetilde{u}\widetilde{v}}$ around $r_{+}$ will be
$g_{\widetilde{u}\widetilde{v}}=\frac{g_{uv}}{\kappa_{+}^{2}(v)}=-W(v)\cdot e^{\kappa_{+}\left(u+v\right)}=-W(v)\cdot e^{\widetilde{u}+\widetilde{v}}$ (6.16)
where $W(v)$ is a non-zero function which depends only on $v$. Transforming into $U,V$ we get:
$g_{UV}=-W(v)$ (6.17)
Our new metric is therefore regular at the EH.
Next we wish to construct initial values for the metric and its derivatives. More specifically, we need the expressions for $R$, $\widetilde{S}$, $R_{,\widetilde{u}}$, $\widetilde{S}_{,\widetilde{u}}$, where $\widetilde{S}$ denotes the value of $S$ in the coordinates $\widetilde{u},\widetilde{v}$, that is, $\frac{e^{\widetilde{S}}}{r}=-g_{\widetilde{u}\widetilde{v}}$. (The derivatives $R_{,\widetilde{u}}$, $\widetilde{S}_{,\widetilde{u}}$ allow us to obtain the initial time derivatives of $R$ and $\widetilde{S}$, see appendix B.) Recall that
$R=r^{2}$ (6.18)
We define:
$r_{*}\equiv\frac{\widetilde{r_{*}}}{\kappa_{+}^{v}(v)}=\frac{\widetilde{u}+\widetilde{v}}{2\kappa_{+}^{v}(v)}$ (6.19)
Therefore, using Eq. (6.7), the derivative of $r$ with respect to $\widetilde{u}$ is:
$\frac{\partial r}{\partial\widetilde{u}}=\left(\frac{\partial r}{\partial r_{*}}\right)_{v}\left(\frac{\partial r_{*}}{\partial\widetilde{u}}\right)_{v}=\frac{f}{2\kappa_{+}^{v}(v)}$ (6.20)
The derivative of $R$ on the initial values is then:
$R_{,\widetilde{u}}=2rr_{,\widetilde{u}}=\frac{rf}{\kappa_{+}(v)}$ (6.21)
To construct $\widetilde{S}$ and its derivative, we first recall that $g_{\widetilde{u}\widetilde{v}}$ on the initial values is:
$\frac{e^{\widetilde{S}}}{r}=-g_{\widetilde{u}\widetilde{v}}=-\frac{f}{2\kappa_{+}(v)^{2}}$ (6.22)
namely,
$\widetilde{S}=\log\left(-\frac{rf}{2\kappa_{+}^{2}(v)}\right)$ (6.23)
The derivative of $\widetilde{S}$ with respect to $\widetilde{u}$ is:
$\widetilde{S}_{,\widetilde{u}}=\left(\frac{\partial}{\partial r}\log\left(-rf\right)\right)r_{,\widetilde{u}}=\left(\frac{\partial}{\partial r}\log\left(-rf\right)\right)\frac{f}{2\kappa_{+}^{2}}$ (6.24)
We got a metric which has a regular geometry, and we found the needed initial values for this metric. However, this metric uses $\widetilde{v},\widetilde{u}$ coordinates which are not convenient to work with. For the numerical simulation, it will be more convenient to use null coordinates similar to Eddington $u,v$ coordinates. In the following subsection, we will move to such more convenient coordinates.
6.3 $g_{1}$ metric
We wish to do coordinate transformation from $\widetilde{g_{1}}$ with $\widetilde{v},\widetilde{u}$ coordinates to new double null coordinates whose scaling (with regards to $M$) is more similar to the Eddington $v,u$ coordinates in RN. (Because it is just coordinate transformation, it will keep the geometry regular at the EH.) For simplicity, we’ll call these new coordinates $v,u$ (however, this new metric should not be confused with the “naive metric” described in subsection 6.1). We shall call the new resulting metric simply “the $g_{1}$ metric”.
The desired coordinate transformation is:
$\widetilde{v}(v)\equiv\mathrel{{\mathop{\int}\limits_{0}^{v}}}\kappa_{+}^{v}(v)dv,\hfill\widetilde{u}(u)\equiv\mathrel{{\mathop{\int}\limits_{0}^{u}}}\kappa_{+}^{u}(u)du$ (6.25)
where $\kappa_{+}^{u}(u)$ was defined in Eq. (4.13). These coordinates satisfy the following relations:
$\frac{d\widetilde{v}}{dv}=\kappa_{+}^{v}(v),\hfill\frac{d\widetilde{u}}{du}=\kappa_{+}^{u}(u)$ (6.26)
Notice that the definition of $\widetilde{u}(u)$ uses $\kappa_{+}^{u}(u)$ rather than $\kappa_{+}^{v}(v)$ because it must be a function of only $u$ (not $v$), so that $u$ will be a null coordinate. The metric function $g_{\widetilde{u}\widetilde{v}}$ now transforms to:
$g_{uv}$ $=g_{\widetilde{u}\widetilde{v}}\frac{\partial\widetilde{u}}{\partial u}\frac{\partial\widetilde{v}}{\partial v}$ (6.27)
$=\left(1-\frac{2M(v)}{r}+\frac{Q(v)^{2}}{r^{2}}\right)\frac{1}{2\kappa_{+}^{v}(v)^{2}}\cdot\left(\kappa_{+}^{v}(v)\kappa_{+}^{u}(u)\right)$ (6.28)
The new line element is therefore:
$ds^{2}=\frac{\kappa_{+}^{u}(u)}{\kappa_{+}^{v}(v)}fdu\;dv+r^{2}d\Omega^{2}$ (6.29)
where, recall,
$f=1-\frac{2M(v)}{r}+\frac{Q(v)^{2}}{r^{2}}$ (6.30)
We’ll now construct the four functions which will be needed for initial values: $R$, $S$, $R_{,u}$, $S_{,u}$. See appendix B for why we require these $u$ derivatives. The function $S$ is:
$S=\log\left(-rg_{uv}\right)=\log\left(-\frac{1}{2}rf\right)+\log\left(\kappa_{+}^{u}(u)\right)-\log\left(\kappa_{+}^{v}(v)\right)$ (6.31)
To get the function $R=r^{2}$, we recall the definition of $r_{*}$ given in Eq. (6.19)
$r_{*}=\frac{\widetilde{u}+\widetilde{v}}{2\kappa_{+}^{v}(v)}$ (6.32)
[…]
$\frac{\widetilde{u}(u)+\widetilde{v}(v)}{2\kappa_{+}^{v}(v)}=r_{*}=r-M+\frac{1}{2\kappa_{+}}\log\left|\frac{r-r_{+}}{M-r_{+}}\right|-\frac{1}{2\kappa_{-}}\log\left|\frac{r-r_{-}}{M-r_{-}}\right|$ (6.33)
The $u$ derivative of $r$ is then (using Eq. (6.20)):
$r_{,u}=r_{,\widetilde{u}}\widetilde{u}_{,u}=f\frac{\kappa_{+}^{u}(u)}{2\kappa_{+}^{v}(v)}$ (6.34)
and therefore:
$R_{,u}=2rr_{,u}=rf\frac{\kappa_{+}^{u}(u)}{\kappa_{+}^{v}(v)}$ (6.35)
The derivative of $S$ with respect to $u$ is:
$S_{,u}=\left(\frac{\partial}{\partial r}\log\left(-rf\right)\right)r_{,u}+\frac{1}{\kappa_{+}^{u}(u)}\frac{d}{du}\kappa_{+}^{u}(u)$ (6.36)
where in $\frac{\partial}{\partial r}\log\left(-rf\right)$ the derivative is taken with $M$,$Q$ being constant:
$\frac{\partial}{\partial r}\log\left(-rf\right)=\frac{1-\frac{Q^{2}}{r^{2}}}{rf}$ (6.37)
Also note that:
$\frac{d}{du}\left(\kappa_{+}^{u}(u)\right)=\frac{\partial}{\partial M}\kappa_{+}^{u}(u)\frac{d}{du}M^{u}(u)+\frac{\partial}{\partial Q}\kappa_{+}^{u}(u)\frac{d}{du}Q^{u}(u)$ (6.38)
Finally, we show that if u+v=0,u+v=0,then $\widetilde{u}+\widetilde{v}=0$ and therefore $r=M(v)$. Choosing a specific value $v=v_{0}$, from $u+v=0$ we obtain that $u=-v_{0}.$ We then get from Eq. (6.25) that:
$\widetilde{v}+\widetilde{u}=\mathrel{{\mathop{\int}\limits_{0}^{v_{0}}}}\kappa_{+}^{v}(v)dv+\mathrel{{\mathop{\int}\limits_{0}^{-v_{0}}}}\kappa_{+}^{u}(u)du$ (6.39)
Using the definition of $\kappa_{+}^{u}(u)$:
$\kappa_{+}^{u}(u)=\kappa_{+}^{v}(v=-u)$ (6.40)
we get that:
$\widetilde{v}+\widetilde{u}$ $=\mathrel{{\mathop{\int}\limits_{0}^{v_{0}}}}\kappa_{+}^{v}(v)dv+\mathrel{{\mathop{\int}\limits_{0}^{-v_{0}}}}\kappa_{+}^{v}(-u)du$ (6.41)
$=\mathrel{{\mathop{\int}\limits_{0}^{v_{0}}}}\kappa_{+}^{v}(v)dv-\mathrel{{\mathop{\int}\limits_{0}^{v_{0}}}}\kappa_{+}^{v}(w)dw=0$ (6.42)
(In the last equality we changed the integration variable w=uw=-u.) We therefore conclude that when $u+v=0$, $\widetilde{u}+\widetilde{v}=0$ and therefore (as seen in subsection 6.2) $r=M(v)$.
6.4 Initial hypersurface and the range of the $v$ and $u$ coordinates
Basically, we have freedom to add constant to the coordinate $v$. Once we’ve chosen this additive constant for $v$, the other coordinate $u$ is uniquely fixed by the condition $r_{*}(r=M)=0$, recalling that $r_{*}\equiv\frac{u+v}{2}$.  [^6] Note that this is the case in the metric $g_{1}$ from which we pick the initial values near the EH. In the semiclassical metric that evolves later on, the relation $r_{*}(r=M)=0$ holds only up to deviations of order $\varepsilon$.
We mark $T\equiv\frac{u+v}{2}$. We choose the initial-value surface to be a surface of constant $T$, which we denote by $T_{0}$. $T_{0}$ is one of the parameters of the numerical simulation, which we set to $T_{0}=-30$ unless stated otherwise. Note that $r-r_{+}=const\cdot e^{2\kappa_{+}T}$, and therefore, the deviation of $r$ at the initial hypersurface from its EH value $r_{+}$ is exponential in $T_{0}$. In all our simulations, $r-r_{+}<4\cdot 10^{-7}$ on the initial hypersurface.
In our numerical simulation, $v$ runs along the initial hypersurface from 0 up to $\Delta v$, where $\Delta v$ is a parameter specifying how large is the numerical domain. We will set $\Delta v=900$ in most cases. Recall that we use units where $M_{0}=1$, where $M_{0}$ is the initial BH mass.
At the point $v=0$ on the initial hypersurface, namely the point marked A in Figure 6.1, $\frac{u+v}{2}=\frac{u}{2}=T_{0}$ and therefore $u=2T_{0}$ there. At the point $v=\Delta v$ on the initial hypersurface, namely the point marked B, $\frac{u+v}{2}=\frac{u+\Delta v}{2}=T_{0}$ and therefore $u=2T_{0}-\Delta v$. Therefore, the choice of the value of $T_{0}$ determines the range of $u$: $u$ runs from $2T_{0}-\Delta v$ to $2T_{0}$. To summarize, the point A is located at $v=0,u=2T_{0}$, the point B at $v=\Delta v,u=2T_{0}-\Delta v$; therefore the point C is at $v=\Delta v,u=2T_{0}-\Delta v$.
The domain of the numerical integration is the grey equilateral right triangle shown in Figure 6.1. The hypotenuse of the triangle is the initial hypersurface, and the two sides are defined by the two null rays $u=2T_{0}$ and $v=\Delta v$.
Figure 6.1: v and u coordinates
We will usually show plots of the various quantities along the two final rays: $u=2T_{0}$ and $v=\Delta v$. The $u=2T_{0}$ ray is called the last u ray, and the $v=\Delta v$ ray is called the last v ray. When we plot quantities along the last $u$ ray $u=2T_{0}$, the plot will be a function of v, and when we plot quantities along the last $v$ ray $v=\Delta v$, the plot will be a function of u. Usually in our simulation, as mentioned above, $\Delta v=900$ and $T_{0}=-30$, so the mentioned two rays are u=60u=-60 and $v=900$.
Part II Numerical results
We now present the results of running the numerical code. Recall that we set our units such that $M_{0}=L_{0}=1$. As a default, we set $du=dv=\frac{1}{100}$, meaning there are hundred grid points per 1 unit of $u$ or $v$. We set the initial-moment parameter $T_{0}=-30$ as a default (unless stated otherwise). Recall that the range of $v$ is usually from $v=0$ to $v=\Delta v=900$, and correspondingly the range of $u$ is from $u=2T_{0}-\Delta v=-960$ to $u=2T_{0}=-60$. Recall that we set $\varepsilon_{0}=\frac{1}{1000}.$
First we will present the results of running with no source terms (which mean that $T_{uv}=T_{\theta\theta}=0$) in section 7. We will present and explore the resulting quantities $R,S$, $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$. In region 2 we will compare the evolving metric versus that of the corresponding local RN metric. Then we turn to explore region 3, namely the near-IH region. We will graphically display both the numerical result and the corresponding analytical approximation from section 4. Then we shall also display the deviation from the analytical approximation caused primarily by finite $\varepsilon_{0}$.
Then we will present the result of running with all sources set to 1 in section 8. We shall again display the quantities of the analytical approximation and their corresponding numerical results, as well as the deviations from the analytical approximation.
Then in the next stage in section 9 we will analyze the consequences of running with finite $T_{0}$ (the value of $T=\frac{u+v}{2}$ on the initial surface) and finite $du$ – by running with several different values of these parameters and comparing the results.
Afterwards in section 10 we will display that the detail of sources doesn’t matter, only the values of $\widetilde{T}_{vv}^{(-)}$ and $\widetilde{T}_{uu}^{(-)}$. We will do it by running with 3 different combinations of the parameters $Z_{R0}$$Z_{Rr}$$Z_{S0}$$Z_{Sr}$ chosen such that they all lead to the same $\widetilde{T}_{vv}^{(-)}$ and $\widetilde{T}_{uu}^{(-)}$. We display the metric functions $R$ and $S$ in these configurations, showing that they are the same in those 3 configurations.
Lastly we display the results of running with three other different combinations of source parameters and show the validity of the analytical approximation in all those runs in section 11
Note that because we are trying to validate the analytical approximation, the numerical simulations do not incorporate the analytical approximation in their assumptions.
7 Numerical results - run with no sources
We start by presenting the results of running with zero source terms, $Z_{R0}=$$Z_{Rr}$$Z_{S0}=$$Z_{Sr}=$0=0, as well as $\delta Q=0$, for $\frac{Q}{M}=0.8$. In this case, $T_{uu}=0$ throughout the whole spacetime, as already mentioned in subsection 2.8 above. We will first look at the entire domain of integration (including all three regions). Then we shall have a more focused look: first on region 2, and then on region 3. In region 2 we will compare the results with the local RN metric. In region 3 we will compare the metric functions to the analytical approximation from section 4. We will then analyze the deviation from the analytical approximation.
7.1 Overview
Figure 7.1: R
(a) $R(v)$ on last u ray
(b) $R(u)$ on last v ray
As one can observe in Figure 7.1a, showing $R(v)$, $R$ starts at a value of $R_{+}\equiv r_{+}^{2}=2.56$ in region 1, drifts down because of the negative $M_{,v}$ (still in region 1); then it “falls down” during the active region (region 2) – as it does in RN – and ends near $R_{-}\equiv r_{-}^{2}=0.16$ near IH, at the beginning of region 3. Note that this apparently rapid decrease in $R$ is actually smooth and regular. This decrease in $R$ from $r_{+}$ to $r_{-}$ happens in a timescale of order a few times $M$ (just like in RN). It only appears sharp because of the horizontal scale which spans a very large range of $v$. Then $R(v)$ drifts slightly further downwards with a small negative $R_{,v}$ according to Eq. (4.40).
The second panel 7.1b, displaying $R(u)$, shows a slightly different behavior. Because $T_{uu}=0$ (unlike $T_{vv}$ which is not zero), $R$ stays almost constant in region 1, $R\approx r_{+}^{2}\left(M(v=900),Q(v=900)\right)=0.551535$, and also in region 3 where $R\approx r_{-}^{2}\left(M(v=900),Q(v=900)\right)=0.034471$. It only varies significantly in region 2, similar to the RN case. In this special case where $J_{u}=0$ and $T_{uu}=0$ the geometry is the charged Vaidya solution, therefore along an incoming null ray with $v=const$ all quantities behave exactly like in the corresponding RN.
Figure 7.2: S
(a) S(v) on last u ray
(b) S(u) on last v ray
The following figure 7.2 shows the behavior of $S$. We first focus on $S(v)$ in panel 7.2a. In region 1, $S_{,v}=\kappa_{+}(v)$, and hence $S$ increases towards region 2 (but this is hard to notice in the figure because of the vertical scale, which mostly shows region 3. The behavior in regions 1 and 2 is better seen in Figure 7.8a below). In region 2, $S$ peaks and then goes down as $S_{,v}=-\kappa_{-}(v)$ on approaching the near-IH region. In region 3, $S$ keeps drifting down over the long stretch from $v=100$ until $v=900$ for the same reason. $S$ eventually reaches very negative values of -4000 and even below. Therefore, $e^{S}\approx 0$ to a very good approximation.
Panel 7.2b is very similar to panel 7.2a, and shows $S$ as a function of $u$, along the other ray $v=900$.
Figure 7.3: $\widetilde{T}_{uu}$ on last v,u rays
(a) $\widetilde{T}_{uu}$ on last u ray
(b) $\widetilde{T}_{uu}$ on last v ray
Next, we show in Figure 7.3 that $\widetilde{T}_{uu}=0$ in the entire domain of integration. The small deviation from zero (primarily seen in region 2) is due to numerical truncation error (we verified this by comparing different $du$ values).
Figure 7.4: $\widetilde{T}_{vv}$ on last u ray
Blue curve is barely seen as it is underneath the orange curve.
We continue to Figure 7.4 which shows both $\widetilde{T}_{vv}$ and the quantity
$2M_{,v}-2\frac{1}{r}QQ_{,v}$ (7.1)
which is the expected value of $\widetilde{T}_{vv}$ according to Eq. (2.107). The two curves overlap, demonstrating that $\widetilde{T}_{vv}$ matches its expected value. In region 2, $\widetilde{T}_{vv}(v)$ goes up from -0.0004 to +0.0004 (due to the rapid change in $r$ in the above expression $2\frac{1}{r}QQ_{,v}$). It then continues to drift in region 3 (though more slowly) as all factors $M_{,v}$ , $r$,$Q$, $Q_{,v}$ appearing in 7.1 slowly drift there with increasing $v$.
Figure 7.5: $\widetilde{T}_{vv}(u)$ on last v ray
Next we see in Figure 7.5 that $\widetilde{T}_{vv}(u)$ is constant in regions 1 and 3, and it only varies in region 2. This is because $M_{,v}$ and $Q_{,v}$ and $Q$ depend only on $v$, hence these quantities are constant along the last $v$ ray. Therefore the only change in the expression 7.1 comes from changing $r$. But $r$ doesn’t change in regions 1 and 3, it only changes in region 2.
Figure 7.6: $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ on initial hypersurface
(a) $\widetilde{T}_{vv}$ on initial hypersurface
(b) $\widetilde{T}_{vv}-\left(2M_{,v}-2\frac{1}{r}QQ_{,v}\right)$ on the initial hypersurface
(c) $\widetilde{T}_{uu}$ on initial hypersurface
In Figure 7.6 we elaborate on the fluxes $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ on the initial hypersurface $T=T_{0}$ (the red line in Figure 6.1). Panel 7.6a compares $\widetilde{T}_{vv}$ along with its theoretically expected value $2M_{,v}-2\frac{1}{r}QQ_{,v}$. $\widetilde{T}_{vv}$ is computed according to 2.42. We display the small difference between these two quantities in Figure 7.6b, which shows the error reaching at most $2\cdot 10^{-5}$. We see that the computed value matches the expected value well. Figure 7.6c, shows that up to numerical errors ($2\cdot 10^{-10}$, mainly roundoff), $T_{uu}=0$ on the initial hypersurface line, as it should.
7.2 Region 2: comparison with the local RN approximation
Now we will show several figures which test the approximation of region 2 as a drifting RN metric.
Figure 7.7: computed R vs. RN value of R
(a) R computed vs. drifting RN on last u ray
(b) R(u) computed vs. RN on last v ray
In Figure 7.7a we compare the numerically-computed $R$ with that corresponding to the RN metric (in double null Eddington coordinates) but with its $M$ and $Q$ parameters replaced with $M(v)$ and $Q^{v}(v)$. Notice the complete overlap of the two curves .
Subsequently, Figure 7.7b shows $R(u)$ along the last $v$ ray, versus RN with constant $M,Q$ corresponding to $M(v=900)$ and $Q^{v}(v=900)$. Notice that because there is no drift of $M$ and $Q$ parameters in the $u$ direction, the two curves nicely overlap.
Figure 7.8: S computed vs. RN
(a) S(v) computed vs. RN on last u ray
(b) S(u) computed vs. RN on last v ray
We move on to show $S$ in Figure 7.8 in region 2, versus the value of $S$ in the corresponding RN metric, determined by
$\frac{e^{S}}{r}=\frac{1}{2}\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)$ (7.2)
In Panel 7.8a we plot $S(v)$ along the u=60u=-60 ray. Then in panel 7.8b we plot $S(u)$ along the last ray $v=900$. In both cases, the two curves are in good agreement. The visible deviation at late times is expected because this range is further from the domain of validity of the local RN approximation which is our focus here.
7.3 Region 3: comparison with the analytical approximation
In this subsection, we will compare the numerically computed quantities $R_{,u},$$R_{,v},$$S_{,u}$$S_{,v}$ in region 3 with their predicted values according to Eqs. (4.5,4.16,4.40,4.41).
Figure 7.9: Numerical result and analytical approximation for $S_{,v}$, $S_{,u}$
(a) Comparison of $S_{,v}$ and $\kappa_{-}^{v}$ on last u ray
(b) Comparison of $S_{,u}$ and $\kappa_{-}^{u}$ on last v ray
We start with Figure 7.9a, which compares $S_{,v}$ and $\kappa_{-}(v)$, both displayed as a function of $v$. They are indeed very close, as predicted by the analytical approximation. Then we see in Figure 7.9b that the approximation is good for $S_{,u}$ as well, when comparing it to $\kappa_{-}^{u}$.
Figure 7.10: Analytical approximation of $R_{,v}$, $R_{,u}$ compared with their numerical values
(a) $R_{,v}$ and $\frac{\widetilde{T}_{vv}^{(-)}}{\kappa_{-}}$ on last u ray
(b) $R_{,u}$ and $\frac{\widetilde{T}_{uu}^{(-)}}{\kappa_{-}}$ on last v ray
Moving on to the derivatives of $R$, we see in Figure 7.10a that the analytical approximation Eq. (4.40) works, comparing $R_{,v}$ with $\frac{\widetilde{T}_{vv}^{(-)}}{\kappa_{-}}$ as they vary as functions of $v$.
Finally, we show in Figure 7.10b the quantities $R_{,u}$ and $\frac{\widetilde{T}_{uu}^{(-)}}{\kappa_{-}^{u}}$ in region 3. These quantities should both vanish, as mentioned in subsection 2.8, because we have no sources in the present run. In the figure we see that both quantities are very small. It is remarkable that their erroneous non-vanishing value is still consistent with Eq. (4.41). Our interpretation of this situation is as follows: these erroneous non-vanishing values happens because of finite-$du$ error. This causes a drift in $\widetilde{T}_{uu}$ leading to non-vanishing $\widetilde{T}_{uu}^{(-)}$– which in turn leads to non-vanishing $R_{,u}$ in region 3, in accordance with Eq. (4.41).
To check this interpretation, we show in Figure 7.11 the magnitude of $R_{,u}$ for two different values of $du$: $\frac{1}{100}$ and $\frac{1}{50}$. The third curve shows the $du=\frac{1}{50}$ case multiplied by $\frac{1}{4}$. We can see that it perfectly overlaps with the curve of $du=\frac{1}{100}$. This shows that the erroneous non-vanishing value of $R_{,u}$ indeed scales with $du^{2}$, implying that it is associated with a finite-$du$ truncation error.
Figure 7.11: $R_{,u}$ with different $du$ values on last v ray
Note that the blue curve completely overlap with the green curve.
7.4 Deviation from the analytical approximation (region 3)
We will now explore and discuss the deviation of the computed evolving $R$ and $S$ from the analytical approximation. We expect these deviations to be caused by finite $\varepsilon_{0}$, as we shall test this expectation below.
Deviation in $S_{,v}$
First, we plot the deviation $S_{,v}+\kappa_{-}$ in Figure 7.12a. (To appreciate the magnitude of this deviation, note that it is roughly about 1% of the value of $S_{,v}$ as seen in Figure 7.9a.)
Figure 7.12: Deviation from analytical approximation of $S_{,v}$, $R_{,v}$
(a) $S_{,v}+\kappa_{-}$ on last u ray
(b) $R_{,v}+\frac{\widetilde{T}_{vv}}{\kappa_{-}}$ on last u ray
We also show the deviation of $R_{,v}$ in Figure 7.12b
To analyze whether this deviation scales with $\varepsilon_{0},$we do two runs: one with $\varepsilon_{0}=\frac{1}{1000}$ as usual, and another one with $\varepsilon_{0}=\frac{1}{500}.$ To compare the two different runs, we plot the mentioned deviations as a function of $M^{v}(v)$ instead of as a function of $v$, to account for the fact that the evaporation rate is different in these two cases. Figure 7.13 compares two different $\varepsilon_{0}$ runs. (Note that in this plot, the low $v$ values correspond to high $M$, with $v=0$ at $M=1$.) If the deviation is proportional to $\varepsilon_{0}$, we would expect it to be twice as big in $\varepsilon_{0}=\frac{1}{500}$ compared to $\varepsilon_{0}=\frac{1}{1000}$.
Figure 7.13: $S_{,v}+\kappa_{-}$ with two $\varepsilon_{0}$ values, scaled by a factor of $\frac{1}{2}$, on the last u ray
To test this assumption, Figure 7.13 scales the curve of $\varepsilon_{0}=\frac{1}{500}$ by $\frac{1}{2}$ (green curve), and now we can see the two curves (green and blue) overlap, confirming that the deviation indeed scales with $\varepsilon_{0}$–meaning that it is a finite-$\varepsilon_{0}$ effect. We also display the unscaled $\varepsilon_{0}=\frac{1}{500}$ curve, as is, in orange. Note that the graph does not start from $M=1$, because here we focus on the analytical approximation of region 3, therefore we restricted the curves to the range of $v$ for which the last $u$ ray is within region 3. In the next figure we see that in earlier $M$ values there is a large deviation from the analytical approximation of region 3.
We also wish to know how this finite-$\varepsilon_{0}$ deviation scales with $M$. From basic semiclassical scaling, one would expect this deviation to scale as $\propto M^{-3}$. This is because the dimensions of $S_{,v}$ is $S_{,v}]=[l]^{-1}$ (as $S$ is dimensionless), and we need to multiply this by the dimensionless factor $\frac{\hbar}{M^{2}}$ (as this deviation in $S_{,v}$ is a semiclassical effect), so the overall scaling goes as $M^{-3}$. We therefore check it by plotting this deviation multiplied by $M^{3},$and check that we get relatively constant line. This approximately constant line is seen in Figure 7.14 (note the narrow range of the vertical scale).
Figure 7.14: $\left(S_{,v}+\kappa_{-}\right)M^{3}$ on last u ray
Deviation in $R_{,v}$
Subsequently, we show in Figure 7.15 the deviation from the analytical approximation for $R_{,v}$ for two different $\varepsilon_{0}$ values. This deviation is equal to $R_{,v}+\frac{\widetilde{T}_{vv}^{(-)}(v)}{\kappa_{-}(v)}$, see Eq. (4.40) .
This time the deviation is expected to scale as $\varepsilon_{0}^{2}$. This is because $R_{,v}$ is proportional to $\varepsilon_{0}$ already at the leading order, so that the finite-$\varepsilon_{0}$ deviation should be proportional to $\varepsilon_{0}^{2}$. We test this assumption by checking that the curve of $\varepsilon_{0}=\frac{1}{500}$ is four times as big as the curve for $\varepsilon_{0}=\frac{1}{1000}$ – by multiplying the curve of $\varepsilon_{0}=\frac{1}{500}$ by a factor of $\frac{1}{4}$.
Figure 7.15: $R_{,v}+\frac{\widetilde{T}_{vv}^{(-)}}{\kappa_{-}}$ with two $\varepsilon_{0}$ values, one of them scaled by $\frac{1}{4}$, on the last u ray
We can see in the figure that the two curves (green and blue) overlap, confirming the scaling by $\varepsilon_{0}^{2}$ as expected.
Figure 7.16: $\left(R_{,v}+\frac{\widetilde{T}_{vv}^{(-)}}{\kappa_{-}}\right)M^{3}$ scaling of the deviation on last u ray
We also wish to know how the mentioned deviation scales with $M$. This deviation scales as $\propto M^{-3}$, and we show it by plotting the deviation times $M^{3},$and check that we get a relatively constant curve. The reason that the deviation scales as $M^{-3}$ is because the dimensions of $R_{,v}]=[l]$ (as $R$ has dimensions of length squared), and we need to further multiply it by a factor of $\varepsilon_{0}^{2}\propto\frac{\hbar^{2}}{M^{4}}$, so that the total power of $M$ becomes $M^{-3}$. This is indeed seen in Figure 7.16 – note the small scale of the vertical axis.
Deviation of $S_{,u}$
The following Figure 7.17 shows the deviation of $S_{,u}$ from $\kappa_{-}^{u}(u)$. Note that we’re now using $M^{u}(u)$ as the horizontal axis, as we need the mass function that depends on $u$. Like the case of $S_{,v}$, we expect the deviation of $\varepsilon_{0}=\frac{1}{500}$ to be twice as big as the deviation for $\varepsilon_{0}=\frac{1}{1000}$.
Figure 7.17: $S_{,u}+\kappa_{-}$ with two $\varepsilon_{0}$ values with one scaled by $\frac{1}{2}$, on the last v ray
We therefore scale the curve of $\varepsilon_{0}=\frac{1}{500}$ by a factor of $\frac{1}{2}$ in Figure 7.17 in green. This figure also shows the unscaled version in orange. The green and blue curves overlap, confirming the scaling by $\varepsilon_{0}$.
[…]
While it might seem that the $\varepsilon_{0}$ deviation is not that small in the present case, note that in astrophysical black holes $\varepsilon_{0}$ is proportional to $10^{-82}$ and so it is a very good approximation for these black holes.
8 Numerical results – run with all sources set to 1
We will now present a numerical run where all the sources are turned on; specifically we shall here choose: $Z_{R0}=$$Z_{Rr}$$Z_{S0}=$$Z_{Sr}=$1=1, and $\alpha=1$ for $\delta Q$ , with $\frac{Q}{M}=0.8$.
8.1 Overview
Figure 8.1: R on last u,v rays
(a) R(v) on last u ray
(b) R(u) on last v ray
As one can observe in Figure 8.1a, showing $R(v)$ along the last-$u$ ray, in region 1 $R$ starts at $r_{+}^{2}(M=1,Q=0.8)=2.56$, and slowly drifts down because of non-vanishing $M_{,v}$. Afterwards it quickly falls down during the active region (region 2) – just like in RN – and ends up near $r_{-}^{2}(M=1,Q=0.8)=0.16$ at the beginning of region 3. Up until region 3, this looks pretty much the same as the case of zero sources shown in Figure 7.1a.
However, in region 3, the two figures look differently: the difference in the value of the sources in these two cases leads to a difference in the corresponding values of $\widetilde{T}_{vv}^{(-)}$. In particular, the two cases have different signs for $\widetilde{T}_{vv}^{(-)}$. Thus, in Figure 8.1a (run with non-vanishing sources) $R(v)$ goes up in region 3 because of negative $\widetilde{T}_{vv}^{(-)}$, whereas in Figure 7.1a (run with vanishing sources) $R$ goes down because of positive $\widetilde{T}_{vv}^{(-)}$. (Recall that $R_{,v}=-\frac{\widetilde{T}_{vv}^{(-)}}{\kappa_{-}}$ and so $R_{,v}$ is of the opposite sign to $\widetilde{T}_{vv}^{(-)}$.)
In the next Figure 8.1b, we look at $R(u)$ along the last $v$ ray. We can see that $R$ starts at $r_{+}^{2}=0.551535$ (corresponding to the final values of $M$ and $Q$ along the EH) and it is almost exactly constant throughout region 1, then it falls down in region 2, similar to Figure 7.1b. Because of the non-vanishing sources, $\widetilde{T}_{uu}$ no longer vanishes. In the present case, $\widetilde{T}_{uu}^{(-)}$ is negative and so $R(u)$ drifts upwards in region 3. This is in contrast to the case of Figure 7.1b where, due to vanishing sources, $\widetilde{T}_{uu}^{(-)}=0$ and $R(u)$ stays constant.
Figure 8.2: $\widetilde{T}_{uu}(v)$ and $\widetilde{T}_{uu}(u)$ on last u,v rays
(a) $\widetilde{T}_{uu}(v)$ on last u ray
(b) $\widetilde{T}_{uu}(u)$ on last v ray
The following Figure 8.2 shows $\widetilde{T}_{uu}$ on the two rays, namely last-$u$ (Figure 8.2a) and last-$v$ (Figure 8.2b) rays. As can be seen in both panels, in region 1 near the initial hypersurface, $\widetilde{T}_{uu}=0$. Then the non-vanishing sources in region 2 cause a non-vanishing, negative $\widetilde{T}_{uu}$.
Figure 8.3: $\widetilde{T}_{vv}(u,v)$
(a) $\widetilde{T}_{vv}(v)$ on last u ray
(b) $\widetilde{T}_{vv}(u)$ on last v ray
In Figure 8.3 we display $\widetilde{T}_{vv}$ on both rays, namely last-$u$ (Figure 8.3a) and last-$v$ (Figure 8.3b) rays. In Figure 8.3a we can see that the chosen source values caused a negative $\widetilde{T}_{vv}$ in region 3, in contrast to Figure 7.4 where $\widetilde{T}_{vv}$ is positive there.
Plotting $\widetilde{T}_{vv}$ as a function of $u$ on the last-$v$ ray in Figure 8.3b, we see that it is almost exactly constant in region 1 and 3, it only changes in region 2 as a result of the non-vanishing sources there.
Figure 8.4: Scaling of $\widetilde{T}_{uu}$ and $\widetilde{T}_{vv}$
(a) $\widetilde{T}_{vv}(v)M^{2}(v)$ on last u ray
(b) $\widetilde{T}_{uu}M^{u}(u)^{2}$ on last v ray
The scaling of both $\widetilde{T}_{vv}^{(+)}$ and $\widetilde{T}_{vv}^{(-)}$ should [^7] This simple scaling holds in our case because $\frac{Q}{M}$ is constant along the EH, and in addition all source terms were constructed with the appropriate $M$ scaling. be $\propto M^{-2}$ and to illustrate that, we plot $\widetilde{T}_{vv}M^{2}$ in Figure 8.4a. This quantity is approximately constant in both region 1 and region 3, confirming that $\widetilde{T}_{vv}\propto M^{-2}$ there.
Next we do a similar check but for $\widetilde{T}_{uu}(u)$ rather than $\widetilde{T}_{vv}(v)$. To this end, we now plot $\widetilde{T}_{uu}M^{u}(u)^{2}$ as a function of $u$ in Figure 8.4b. This quantity is again approximately constant in region 1 and region 3, confirming that $\widetilde{T}_{uu}\propto M^{u}(u)^{-2}$ in both these regions.
Because the sources only affect the metric during and after region 2, in region 1 the results are (almost) exactly the same as the no-sources run. In particular the behavior of $\widetilde{T}_{vv}(v)$ on the initial hypersurface in the present case is exactly the same as seen in Figure 7.6a.
8.2 Region 3: comparison with the analytical approximation
Like in the corresponding subsection 7.3, we will now compare $R_{,u},$$R_{,v},$$S_{,u}$$S_{,v}$ with their expected values according to Eqs. (4.5,4.16,4.40,4.41).
Figure 8.5: Analytical approximation of $S_{,v}$ and $S_{,u}$
(a) $S_{,v}$ and $\kappa_{-}$ on last u ray
(b) $S_{,u}$ and $\kappa_{-}^{u}$ on last v ray
The first Figure 8.5a, show $S_{,v}(v)$ and $\kappa_{-}(v)$ along the last-$u$ ray. They are indeed very close, as predicted by the analytical approximation. However, there is a small visible gap near the end at around $v\sim 900$ (which is almost invisible in the corresponding Figure 7.9a). We will show in the next subsection that this is a finite-$\varepsilon_{0}$ deviation.
We see in the following Figure 8.5b that the approximation is good for $S_{,u}$ as well. Similar to the $S_{,v}$case, there is a small visible gap near the end at u=950u=-950 (which doesn’t appear at the corresponding Figure 7.9b). This is again a finite-$\varepsilon_{0}$ deviation.
Figure 8.6: Analytical approximation of $R_{,v}$ and $R_{,u}$
(a) $R_{,v}$ and $\frac{\widetilde{T}_{vv}}{\kappa_{-}}$ on last u ray
(b) $R_{,u}$ and $\frac{\widetilde{T}_{uu}}{\kappa_{-}^{u}}$ on last v ray
We move on to explore the analytical approximation for $R$. In Figure 8.6a, we see that the analytical approximation for $R_{,v}$ Eq. (4.40) works well. In this case, in contrast to Figure 7.10a, $R_{,v}>0$ because in the present case the chosen sources lead to $\widetilde{T}_{vv}<0$.
Finally, we show in Figure 8.6b that the analytical approximation for $R_{,u}$ Eq. (4.41) works as well. This time, unlike Figure 7.10b (the case of vanishing sources), $\widetilde{T}_{uu}\neq 0$ because of the non-vanishing sources.
8.3 Impact of $\delta Q$
In this run we used $\alpha=1$ parameter for $\delta Q$ (unlike the previous run which had $\alpha=0$). First we will show in Figure 8.7 the value of $Q(u,v)$ along the last-$u$ ray (including the $\delta Q$ term) and compare it with $Q^{v}(v)$ which is the value of $Q$ at the EH. As seen in the figure, the value of $Q$ drifts by a lot and reaches 4 at maximal $v$, meaning that $Q(u,v)$ is greater than $M(v)$ – which might naively be regarded as “overcharging”.
Figure 8.7: $Q^{v}$ and $Q=Q^{v}+\delta Q$ on the last u ray
Recall that the analytical approximation for $S_{,v}$ is $\kappa_{-}^{v}$, and that $\kappa_{-}^{v}(v)$ was defined to be $\kappa_{-}^{v}(M(v),Q^{v}(v))$, see Eq. (4.6). One might be tempted to think that perhaps the more relevant charge parameter, to be used in $\kappa_{-}^{v}(v)$, would be $Q(u,v)=Q^{v}(v)+\delta Q(u,v)$ rather than just $Q^{v}(v)$. The purpose of the next figure is to check this issue.
In Figure 8.8 we show $S_{,v}$ along with two “candidate” $\kappa_{-}^{v}$ curves: the green with the correct choice $Q^{v}(v)$, and red with the incorrect choice $Q(u,v)=\delta Q+Q^{v}$. As can be clearly seen in the figure, the green curve matches the blue curve, while the red curve quickly deviates from them in region 3. (The red curve is cut short because $M$ becomes bigger than $Q(u,v)$, meaning the corresponding $\kappa_{-}$ would have a negative square root.) This figure indicates that using $Q^{v}(v)$ in the analytical approximation is indeed the correct choice. This is not surprising, of course, because our construction of the analytical approximation for $S_{,v}$ in subsection 4.3 provided the $S_{,v}=-\kappa_{-}^{v}$ specifically with the charge parameter $Q^{v}(v)$, see Eq. (4.6).
Figure 8.8: $S_{,v}$ and $\kappa_{-}$ on last u ray, using both $Q^{v}$ and $Q=\delta Q+Q^{v}$ for $\kappa_{-}$
8.4 Deviations from the analytical approximation
In this section we intend to analyze the deviation from the analytical approximation in order to test whether it is a finite-$\varepsilon_{0}$ effect. To this end, similar to subsection 7.4 above, we’ll compare the deviation from the analytical approximation for two different $\varepsilon_{0}$ values: $\frac{1}{500}$ and $\frac{1}{1000}$. We will again scale the $\varepsilon_{0}=\frac{1}{500}$ curve by a factor of $\frac{1}{2}$ for $S$ and $\frac{1}{4}$ for $R$. As before, we will plot these scaled deviations as a function of the mass $M(v)$. We will see that the two curves overlap, confirming that the deviation scales as $\varepsilon_{0}$ for $S$ and as $\varepsilon_{0}^{2}$ for $R$. We shall also show the unscaled $\varepsilon_{0}=\frac{1}{500}$ curve in orange.
Figure 8.9: $S_{,v}+\kappa_{-}$ with two $\varepsilon_{0}$, scaled by $\frac{1}{2}$ on last u ray
The first Figure 8.9 shows the deviation of $S_{,v}$ for the two mentioned values of $\varepsilon_{0}$, with the $\varepsilon_{0}=\frac{1}{500}$ deviations scaled by $\frac{1}{2}$. The two curves (green and blue) indeed overlap, confirming the scaling of the deviation as $\varepsilon_{0}.$ Comparing to Figure 7.13 (vanishing sources run), we see that the deviation is larger by around a factor of $\sim 2$ in the current case, but this change is of no importance.
Figure 8.10: $S_{,u}+\kappa_{-}^{u}$ with two $\varepsilon_{0}$, scaled by $\frac{1}{2}$ on last v ray
The second Figure 8.10 shows the deviation of $S_{,u}$ for the same two values of $\varepsilon_{0}$, with the $\varepsilon_{0}=\frac{1}{500}$ deviations scaled by $\frac{1}{2}$. The two curves (blue and green) again overlap, confirming a scaling of the deviation in $S_{,u}$ as $\varepsilon_{0}$.
Figure 8.11: $R_{,v}+\frac{\widetilde{T}_{vv}}{\kappa_{-}}$ with two $\varepsilon_{0}$ and scaled by $\frac{1}{4}$ on last u ray
The following Figure 8.11 shows the deviation of $R_{,v}$ for the two values of $\varepsilon_{0}$, with $\varepsilon_{0}=\frac{1}{500}$ curve scaled by $\frac{1}{4}$. The two curves overlap, confirming scaling with $\varepsilon_{0}^{2}$ – as expected from $R$ derivatives in region 3, see explanation in subsection 7.4.
Finally we explore the deviation of $R_{,u}$ from its analytical approximation. Unlike the no-sources run considered in subsection 7.4, this time there is non-zero $\widetilde{T}_{uu}$ and hence non-vanishing $R_{,u}$ in region 3, and Figure 8.12 shows the deviation of $R_{,u}$, with the $\varepsilon_{0}=\frac{1}{500}$ curve scaled by $\frac{1}{4}$. The two curves (green and blue) again overlap, confirming the scaling of the deviation as $\text{$\varepsilon_{0}^{2}$}.$ (We expect the same scaling of the deviation for $R_{,u}$ as for $R_{,v}$.)
Figure 8.12: $R_{,u}+\frac{\widetilde{T}_{uu}}{\kappa_{-}^{u}}$ with two $\varepsilon_{0}$ and scaled by $\frac{1}{4}$ on last v ray
9 Finite-$du$ and finite-$T_{0}$ errors
In this section, we shall explore two kinds of errors:
  1. 1.
    The numerical error associated with finite $du$ (namely the truncation error),
  2. 2.
    The error associated with setting the initial values on a hypersurface of finite $T_{0}$ value – rather than at the event horizon itself, which would correspond to $T_{0}\to$-$\infty$.
Specifically, we will work here with the set of parameters of section 8:
$Z_{R0}=$$Z_{Rr}$$Z_{S0}=$$Z_{Sr}=$$Z_{*}=1$, and $\alpha=1$ for $\delta Q$, and with $\frac{Q}{M}=0.8$. We shall focus primarily on the magnitude of these errors in region 3.
9.1 Finite-$du$ error
Our algorithm discretizes spacetime into a two-dimensional grid with separation $du$ (we use $du=dv$) between two adjacent grid locations. Because the distance between grid points is finite, we have an error in our numerical calculation, which increases with the size of $du$. We would like to estimate this error which results from having finite $du$.
To estimate the magnitude of this error, we run the numerical code with three different $du$ values: $du=\frac{1}{50}$, $du=\frac{1}{100}$, and $du=\frac{1}{200}$, and compare the results. The difference between the runs $du=\frac{1}{50}$ and $du=\frac{1}{100}$ provides an estimate for the truncation error in the $du=\frac{1}{50}$ run (up to a factor $\frac{4}{3}$ which we explain below), and similarly the difference between the $du=\frac{1}{100}$ and $du=\frac{1}{200}$ runs provides an estimate for the truncation error in the $du=\frac{1}{100}$ run, as we will now explain.
Let $f\left(du\right)$ denote some numerically computed quantity (that necessarily depends on $du$), and let $e\left(du\right)$ denote the numerical error in the computation of that quantity. We expect the error to scale as $du^{2}$, as already mentioned above in section 5 (and as will be confirmed below):
$e\left(du\right)=\beta\cdot du^{2}$ (9.1)
where $\beta$ is a certain parameter. Under this expectation, we evaluate the difference between the value of the quantity $f$ computed with two different $du$ values:
$f\left(du\right)-f\left(\frac{1}{2}du\right)$ $=e\left(du\right)-e\left(\frac{1}{2}du\right)$ (9.2)
$=\beta\left(du^{2}-\left(\frac{1}{2}du\right)^{2}\right)=\frac{3}{4}\beta du^{2}$ (9.3)
Therefore, we can compute the parameter $\beta$:
$\beta=\frac{4}{3}\frac{f\left(du\right)-f\left(\frac{1}{2}du\right)}{du^{2}}$ (9.4)
Thus, the error in $du$ is equal to:
$e\left(du\right)=\frac{4}{3}\left(f\left(du\right)-f\left(\frac{1}{2}du\right)\right)$ (9.5)
In particular, if we wish to estimate the numerical error $e\left(\frac{1}{100}\right)$ given the values of a certain quantity $f$ at $du=\frac{1}{100}$ and $du=\frac{1}{200}$, we get:
$e\left(\frac{1}{100}\right)=\frac{4}{3}\left[f\left(\frac{1}{100}\right)-f\left(\frac{1}{200}\right)\right]$ (9.6)
(and similarly for $du=\frac{1}{50}$).
Figure 9.1: Truncation error in $R$ on last u ray
We first present the results for the truncation error in $R$ in Figure 9.1, which shows the two differences between running with different $du$ values. Note how in region 1 the error is very close to zero. The error for $du=\frac{1}{100}$ is of order $\sim 10^{-5}$ as indicated by the orange curve.
Figure 9.2: Ratio of truncation errors $\frac{R\left[du=\frac{1}{50}\right]-R\left[du=\frac{1}{100}\right]}{R\left[du=\frac{1}{100}\right]-R\left[du=\frac{1}{200}\right]}$ on last u ray
We next present the ratio of these two differences in Figure 9.2. As expected, the ratio is equal to 4, confirming error scaling as $du^{2}$. We point out that the error in region 1 (not seen in this figure) is exceptional: it is very small (see Figure 9.1) and it doesn’t scale as $du^{2}$. In this region, the truncation error is negligible and other errors seem to dominate.
We shall now present similar findings for other quantities $S$, $S_{,v}$$R_{,v}$:
Figure 9.3: Truncation error in S on last u ray
Figure 9.3 shows the truncation error in $S$. Note that while the error reaches a relatively high value, 0.04, $S$ itself reaches $\sim 4000$ so the relative error is again of order $10^{-5}$.
Figure 9.4: Ratio of truncation errors $\frac{S\left[du=\frac{1}{50}\right]-S\left[du=\frac{1}{100}\right]}{S\left[du=\frac{1}{100}\right]-S\left[du=\frac{1}{200}\right]}$ on last u ray
We display the ratio of these two truncation errors in Figure 9.4, and it can be seen that in region 2 and 3 the ratio is 4, confirming error scaling with $du^{2},$as before. In region 1, the error is very small and doesn’t scale with $du$ for the same reason explained above.
Figure 9.5: Truncation error in $R_{,v}$ on last u ray
Next we show the errors in $R_{,v}$ in Figure 9.5. In region 1, the errors are roughly equal because truncation error isn’t the leading cause of errors. In region 3 they seem to scale as expected. We computed the ratio in Figure 9.6, and it appears concentrated around 4, as expected from $du^{2}$ error scaling. It appears more jittery because of round-off error.
Figure 9.6: Ratio of truncation errors $\frac{R_{,v}\left[du=\frac{1}{50}\right]-R_{,v}\left[du=\frac{1}{100}\right]}{R_{,v}\left[du=\frac{1}{100}\right]-R_{,v}\left[du=\frac{1}{200}\right]}$ on last u ray
Figure 9.7: Truncation error in $S_{,v}$ on last u ray
In Figure 9.7 we show the error in $S_{,v}$. It is of order $10^{-4}.$ The error in region 1 is again close to zero.
Figure 9.8: Ratio of truncation errors $\frac{S_{,v}\left[du=\frac{1}{50}\right]-S_{,v}\left[du=\frac{1}{100}\right]}{S_{,v}\left[du=\frac{1}{100}\right]-S_{,v}\left[du=\frac{1}{200}\right]}$ on last u ray
We see in Figure 9.8 that the ratio is also around 4 in region 3, as expected from $du^{2}$ scaling.
We shall not show the truncation errors in $R_{,u}$ and $S_{,u}$ because they are similar to the corresponding truncation errors in $R_{,v}$ and $S_{,v}$.
Based on all these tests, we conclude that the truncation error in our numerical code indeed scales as $du^{2}$ (as it should).
9.2 Finite $T_{0}$ errors
Ideally, we would choose $T_{0}=$-$\infty$ and start the evolution at the event horizon. However, we can’t do that because that would require infinite precision (as the difference of $R$ from $r_{+}^{2}$ is exponentially small in $T_{0}$); and furthermore, at a certain point choosing too negative $T_{0}$ value would result in increasing errors. Our goal here is to quantify the error resulting from choosing a finite value for $T_{0}$. To do that, we run the code with four different values of $T_{0}$, and analyze the results, as explained below. We compare the results of the different runs on a common $u=const$ ray – namely on a certain $u=const$ ray that appears in the numerical grids of all 4 runs. (Note that such a common $u=const$ ray corresponds to different grid indices in the different runs.)
Specifically, we run the code with the following values of $T_{0}$: 20,25,30,35-20,-25,-30,-35. In principle we expect the finite-$T_{0}$ error to be exponentially small in $T_{0}$. Therefore we expect this error to be very small for all 4 cases. Yet the differences are measurable (see below) and we shall analyze these differences. We chose the common ray to be u=80u=-80 (which is close to the maximal $u$ values in all these grids).
Figure 9.9: Difference in R between different $T_{0}$ values on last u ray
In figure 9.9, we show the difference in $R$ between the three pairs of successive $T_{0}$ runs. We see in the graph that this difference decreases by a factor of approximately 10 between two such successive pairs. Notice that $e^{2\kappa_{+}\Delta T_{0}}\approx 10.4$ where $\Delta T_{0}=5$ is the difference in $T_{0}$ between two successive runs, which nicely agrees with this factor of $\sim 10$.
Throughout this work we use, as default, $T_{0}=-30$, and the corresponding finite $T_{0}$ error is estimated to be approximately of order $10^{-8}$ in region 3 (as seen in the green curve).
Figure 9.10: Difference in S between different $T_{0}$ values on last u ray
In figure 9.10, we show the difference in $S$ between the three pairs of successive $T_{0}$ runs. We can see a similar behavior for $S$ as it was for $R$: the gaps here again correspond to the error being reduced by a factor of approximately 10. Here the error in $T_{0}=-30$ is approximately of order $10^{-5}$.
Figure 9.11: Difference in $R_{,v}$ between different $T_{0}$ values on last u ray
In the following Figure 9.11, we see the differences in $R_{,v}$ between pairs of successive $T_{0}$ runs. In the region $v>600$ we see that the round-off error starts to dominate over the finite-$T_{0}$ error. The error in $T_{0}=-30$ is found to be of order $10^{-10}$ (or smaller) in region 3.
Figure 9.12: Difference in $S_{,v}$ between different $T_{0}$ values on last u ray
Moving on to Figure 9.12, which shows the corresponding differences in the values of $S_{,v}$, we again see that the roundoff error dominates at $v>600$. The finite-$T_{0}$ error in $T_{0}=-30$ is of order $10^{-7}$.
$S_{,u}$ and $R_{,u}$ display similar features to $S_{,v}$ and $R_{,v}$ respectively: the finite-$T_{0}$ error is reduced by a factor of $\sim 10$ when $T_{0}$ decreases by 5. We shall not display them here.
In conclusion, for our standard choice $T_{0}=-30$, all the above quantities show that the finite-$T_{0}$ error is negligibly small.
10 Independence from the detail of sources
As was already mentioned in subsection 4.8, our analytical approximation for region 3 only depends on the four input functions of $M(v)$, $Q(v)$, $\widetilde{T}_{vv}^{(-)}(v)$ and $\widetilde{T}_{uu}^{(-)}(v)$. Once $M(v)$ and $Q(v)$ are prescribed, they directly determine $\widetilde{T}_{vv}^{(+)}$, and $\widetilde{T}_{uu}^{(+)}=0$ from regularity. Then $\widetilde{T}_{vv}^{(-)}$ and $\widetilde{T}_{uu}^{(-)}$ are in principle determined by the evolution from the EH to the near-IH region (with $\widetilde{T}_{vv}^{(+)}$ and $\widetilde{T}_{uu}^{(+)}=0$ serving as initial values) – in a manner that depends on the sources $Z_{R}$ and $Z_{S}$. Thus, these sources do affect $\widetilde{T}_{vv}^{(-)}$ and $\widetilde{T}_{uu}^{(-)}$ – but otherwise the analytical approximation is independent of these sources. Our goal in this section is to numerically check this independence from the details of sources.
To check this independence, we run the numerical code with three different sets of sources which all lead to (approximately) the same values of $\widetilde{T}_{vv}^{(-)}$ and $\widetilde{T}_{uu}^{(-)}$, and we show that the results for $R_{,v}$$R_{,u}$$S_{,v}$$S_{,u}$ are the same in all three cases – meaning they are independent of the specific choice of the sources. We use the parametrization given in Eqs. (3.23,3.24) for $Z_{R}$ and $Z_{S}$ in terms of the four parameters $Z_{R0}$, $Z_{Rr}$, $Z_{S0}$, $Z_{Sr}$. In all three runs, we choose $\alpha=1$ for $\delta Q$, and $\frac{Q}{M}=0.8$. The three sets of source parameters used in these three runs are given in Table 1.
Table 1: Runs with equal $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$
run $Z_{R0}$ $Z_{Rr}$ $Z_{S0}$ $Z_{Sr}$ $Z_{S0}+Z_{Sr}-\frac{1}{2}Z_{Rr}$
run1 1 1 1 1 1.5
run2 1 -1 0 1 1.5
run3 1 1 0 2 1.5
We found empirically that what matters for the final values of $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ in region 3 is the combination $Z_{S0}+Z_{Sr}-\frac{1}{2}Z_{Rr}$, as displayed at the last column. (This empirical expression may be explored analytically, but this is beyond the scope of this research.) In the first stage we shall check that in region 3 we have the same $\widetilde{T}_{uu}$ and $\widetilde{T}_{vv}$ for all three sets of parameters.
Figure 10.1: $\widetilde{T}_{uu}$ for the three different runs on the last u ray
Figure 10.1 shows that $\widetilde{T}_{uu}$ is approximately equal for the different runs (the curves overlap).
Figure 10.2: $\widetilde{T}_{uu}$ for the three different runs, in region 3, on the last u ray
In the next Figure 10.2 we zoom in to region 3 where we compare $\widetilde{T}_{uu}$ for the three different runs, and we can see they are pretty close to each other, with difference of order $\sim 10^{-5}$, to be compared with the $\sim 10^{-3}$ value of $\widetilde{T}_{uu}$.
Figure 10.3: $\widetilde{T}_{vv}$ for the three different runs on the last u ray
In the following Figure 10.3, we show that $\widetilde{T}_{vv}$ is also approximately equal.
Figure 10.4: $S_{,v}$ for the three different runs, compared to $\kappa_{-}(v)$ on the last u ray
We move on to show $S_{,v}$ along with its expected value $\kappa_{-}(v)$ in Figure 10.4. The analytical approximation works well for all runs.
Figure 10.5: $S_{,v}$ for the three different runs, compared to $\kappa_{-}(v)$, zoomed in on the last u ray
We zoom in to the region $800\leq v\leq 900$ in Figure 10.5 to show how close are the three different runs.
Figure 10.6: $S_{,u}$ for the three different runs, compared to $\kappa_{-}^{u}$ on the last v ray
We continue to show $S_{,u}$ and its expected value $\kappa_{-}^{u}(u)$ in Figure 10.6. All the three equivalent runs are close together in region 3, near their expected value of $\kappa_{-}^{u}(u)$.
Figure 10.7: $R_{,v}$ for the three different runs, compared to $\frac{\widetilde{T}_{vv}}{\kappa_{-}^{v}}$ on the last u ray
We turn now to explore the derivatives of $R$. Figure 10.7 shows both $R_{,v}$ and its analytical approximation $\frac{\widetilde{T}_{vv}}{\kappa_{-}^{v}}$ for each run. All the three runs and their approximations are close together.
Figure 10.8: $R_{,u}$ for the three different runs, compared to $\frac{\widetilde{T}_{uu}}{\kappa_{-}^{u}}$ on the last v ray
Subsequently, we show in Figure 10.8 both $R_{,u}$ and its analytical approximation $\frac{\widetilde{T}_{uu}}{\kappa_{-}^{u}}$ for each run. All curves are close together, showing that the approximation is good.
In conclusion, we have demonstrated that for all quantities we explored here ($R_{,v}$$R_{,u}$$S_{,v}$$S_{,u}$), the three different runs indeed yielded approximately the same results regardless of which specific sources were used – and that the corresponding analytical approximation is good in all cases.
11 Three additional runs with different sources
We shall now extend the validity test of the analytical approximation to three additional sets of source parameters – this time we chose parameter sets which lead to different values of $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ in region 3 (unlike what we did in section 10). Table 2 details the various source parameters ($Z_{R0}$, $Z_{Rr}$, $Z_{S0}$, $Z_{Sr}$$\alpha$) and $\frac{Q}{M}$ used in each run. Run0 was detailed in section 7. Run1 was detailed in section 8. Our focus here is runs 2 through 4.
In all three runs we shall display $R(v)$,$R(u)$, and then the fluxes $\widetilde{T}_{vv}(v)$ and $\widetilde{T}_{uu}(u)$, followed by the 4 derivatives $S_{,v}(v)$$S_{,u}(u)$ and $R_{,v}(v)$$R_{,u}(u)$. All these quantities will be displayed along either the last $u=const$ or $v=const$ rays depending on their arguments ($u$ or $v$). As we shall see in all these 3 runs, the analytical approximation works well.
Table 2: Runs with equal $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$
The signs of $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ are shown
run $Z_{R0}$ $Z_{Rr}$ $Z_{S0}$ $Z_{Sr}$ $\alpha$ $\frac{Q}{M}$ $Z_{S0}+Z_{Sr}-\frac{1}{2}Z_{Rr}$ $\widetilde{T}_{vv}$ $\widetilde{T}_{uu}$
run0 0 0 0 0 0 0.8 0 + 0
run1 1 1 1 1 1 0.8 1.5
run2 1 1 -1 -1 $\frac{1}{2}$ 0.8 -2.5 +
run3 -1 -1 1 1 1 0.6 2.5
run4 1 $\frac{1}{2}$ 1 $\frac{1}{2}$ $\frac{1}{2}$ 0.8 0.75 + +
11.1 Run2
This run is done with sources equal to $Z_{R0}=$$Z_{Rr}=1$, $Z_{S0}$$Z_{Sr}=$1=-1, and $\alpha=\frac{1}{2}$ for $\delta Q$ and with $\frac{Q}{M}=0.8$.
We can see in Figure 11.2a that $\widetilde{T}_{vv}>0$ in region 3, and from Figure 11.2b $\widetilde{T}_{uu}<0$ in region 3. This explains why in Figure 11.1a $R$ is decreasing with $v$ ($R_{,v}<0$) in region 3, while in Figure 11.1b $R$ is increasing with $u$ ($R_{,u}>0$) in region 3. Figure 11.3a shows the value of $S_{,v}$ is close to its analytical approximation. Similarly, in Figure 11.3b we can see that $S_{,u}$ is also close to $\kappa_{-}^{u}$, as expected from the analytical approximation. Figure 11.4a compares the value of $R_{,v}$ to the analytical approximation, and demonstrates that they are close. Figure 11.4b compares the value of $R_{,u}$ to the analytical approximation, showing they are close too.
Figure 11.1: R for run2
(a) $R(v)$ for run2 on last u ray
(b) $R(u)$ for run2 on last v ray
Figure 11.2: $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ for run2
(a) $\widetilde{T}_{vv}$ for run2 on last u ray
(b) $\widetilde{T}_{uu}$ for run2 on last v ray
Figure 11.3: $S_{,v}$ and $S_{,u}$ and their analytical approximation for run2
(a) $S_{,v}$ and $\kappa_{-}$ for run2 on last u ray
(b) $S_{,u}$ and $\kappa_{-}^{u}$ for run2 on last v ray
Figure 11.4: $R_{,v}$ and $R_{,u}$ and their analytical approximation for run2
(a) $R_{,v}$ and $\frac{\widetilde{T}_{vv}}{\kappa_{-}}$ for run2 in region 3 on last u ray
(b) $R_{,u}$ and $\frac{\widetilde{T}_{uu}}{\kappa_{-}^{u}}$ for run2 in region 3 on last v ray
11.2 Run3
This run is done with sources equal to $Z_{R0}=$$Z_{Rr}=-1$, $Z_{S0}$$Z_{Sr}=$1=1, and $\alpha=1$ for $\delta Q$. This time (and unlike all other cases) we choose $\frac{Q}{M}=0.6$.
As can be seen in Figure 11.5a, $R$ reaches its minimum value at the end of region 2 and then drifts upwards in region 3. This minimal value of $R=r_{-}^{2}=0.04$ is quite small (compared to e.g. in Figure 11.1a) for a simple reason: Because $\frac{Q}{M}$ is now smaller, $r_{-}$ is smaller than in previous runs when $\frac{Q}{M}$ was chosen to be 0.8. Figures 11.6a and 11.6b respectively show that both $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ are <0<0. Because of that, $R_{,u}>0$ and also $R_{,v}>0$ in region 3, as can be seen by the increasing $R$ in Figures 11.5a, 11.5b.Figures 11.7,11.8 show the quantities $S_{,v}$$S_{,u}$ and $R_{,v}$$R_{,u}$ respectively, compared with their analytical approximation values. It can be seen that the analytical approximation is close to the numerically computed values for all these quantities.
Figure 11.5: R for run3
(a) $R(v)$ for run3 on last u ray
(b) $R(u)$ for run3 on last v ray
Figure 11.6: $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ for run3
(a) $\widetilde{T}_{vv}$ for run3 on last u ray
(b) $\widetilde{T}_{uu}$ for run3 on last v ray
Figure 11.7: $S_{,v}$ and $S_{,u}$ and their analytical approximation for run 3
(a) $S_{,v}$ and $\kappa_{-}$ for run3 on last u ray
(b) $S_{,u}$ and $\kappa_{-}^{u}$ for run3 on last v ray
Figure 11.8: $R_{,v}$ and $R_{,u}$ and their analytical approximation for run3
(a) $R_{,v}$ and $\frac{\widetilde{T}_{vv}}{\kappa_{-}}$ for run3 on last u ray
(b) $R_{,u}$ and $\frac{\widetilde{T}_{uu}}{\kappa_{-}^{u}}$ for run3 on last v ray
11.3 Run4
This run is done with source parameters taken to be $Z_{R0}=Z_{S0}$=1=1, $Z_{Rr}$$Z_{Sr}=$$\frac{1}{2}$, and $\alpha=-\frac{1}{2}$ for $\delta Q$ and with $\frac{Q}{M}=0.8$. In this run, unlike all other runs, the maximum $v$ is taken to be 482 (compared with 900 in all previous runs) to avoid (or at least reduce) complications that may arise from the singularity that is going to form in this case (see below) – except for Figure 11.10b where $v$ is taken to be 600 in order to show where the singularity forms on this null ray.
Figures 11.11a and 11.11b respectively show that this time, both $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ are >0>0 in region 3. Because of that, both $R_{,v}$ and $R_{,u}$ are negative there. Correspondingly, Figure 11.9 shows that, at u=60u=-60, $R$ is decreasing and reaching zero at $v=482$. Likewise, Figure 11.10b shows, at $v=600$, that $R$ is also decreasing as a function of $u$, reaching zero at u=274u=-274, which is marked by a short red line. This vanishing of $R$ – seen in all these figures – expresses the formation of a spacelike $R=0$ singularity, which we shall further discuss at the end of this subsection. We also show Figure 11.10a, which shows that at $v=482$, $R$ is decreasing all the way to zero at u=60u=-60. We chose to show quantities on this ray of $v=482$ for the rest of this case to avoid crossing the singularity, as this is the value of $v$ at which $R$ vanishes along the last-$u$ ray u=60u=-60. Figure 11.12a and 11.12b respectively show $S_{,v}$ and $S_{,u}$ versus their analytical approximations $\kappa_{-}(v)$ and $\kappa_{-}^{u}(u)$, again showing the good quality of the analytical approximation. Figure 11.13a shows $R_{,v}$ compared with its analytical approximation, and they are close but have a small gap which is this time bigger relative to the previous runs. A similar gap is also seen in Figure 11.13b which shows $R_{,u}$ compared with its analytical approximation. We shall now explore this gap, and show that this deviation from the analytical approximation is a finite-$\varepsilon_{0}$ deviation (which in this specific case happens to be bigger than in the previous cases).
Figure 11.9: $R(v)$ for run4 on last u ray
Figure 11.10: $R(u)$ for run4 along two $v=const$ rays
(a) $R(v=482,u)$ for run4
(b) $R(v=600,u)$ for run4
The vertical red line in Panel b denotes the $u$-value where $R=0$
Figure 11.11: $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ for run4
(a) $\widetilde{T}_{vv}$ for run4 on last u ray
(b) $\widetilde{T}_{uu}$ for run4 on last v ray
Figure 11.12: $S_{,v}$ and $S_{,u}$ and their analytical approximation for run4
(a) $S_{,v}$ and $\kappa_{-}$ for run4 on last u ray
(b) $S_{,u}$ and $\kappa_{-}^{u}$ for run4 on last v ray
Figure 11.13: $R_{,v}$ and $R_{,u}$ and their analytical approximation for run4
(a) $R_{,v}$ and $\frac{\widetilde{T}_{vv}}{\kappa_{-}}$ for run4 on last u ray
(b) $R_{,u}$ and $\frac{\widetilde{T}_{uu}}{\kappa_{-}^{u}}$ for run4 on last v ray
To this end, we shall run the same case with $\varepsilon_{0}=\frac{1}{500}$ as well, and compare the deviations from the analytical approximation obtained in both values of $\varepsilon_{0}$, namely $\frac{1}{500}$ and $\frac{1}{1000}$, similar to subsection 7.4. Figure 11.14 presents this deviation in $R_{,v}$ for these two values of $\varepsilon_{0}$, with the $\varepsilon_{0}=\frac{1}{500}$ case scaled by $\frac{1}{4}$. The blue and orange curves overlap, demonstrating that this deviation from the analytical approximation scales as $\varepsilon_{0}^{2}$ – indicating that this is indeed a finite-$\varepsilon_{0}$ deviation.
Figure 11.14: $R_{,v}+\frac{\widetilde{T}_{vv}}{\kappa_{-}}$ with two $\varepsilon_{0}$ and scaled by $\frac{1}{4}$ in run4 on last u ray
Figure 11.15: Spacetime diagram for $R=0$ singularity
In this case (and unlike the previous cases we explored above), because both $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ are positive, a spacelike $R=0$ singularity forms at the middle of the numerical domain, blocking access to the would-be CH. The spacetime diagram corresponding to this situation is shown in Figure 11.15.This $r=0$ spacelike singularity that forms inside the semiclassical charged black hole resembles the spacelike singularity found in Ref [6], although it must be recalled that they studied classical perturbations whereas here we consider the backreaction of a quantum field.  [^8] There is a possibility that the two fluxes will have opposite signs in region 3. In this case, as it turns out, what matters is the sign of $\widetilde{T}_{uu}$: if it is negative (as it is e.g. in run 2 above), then a CH will form, and if it is positive, a spacelike $R=0$ singularity will form instead.
12 Conclusion
The main research question that motivated us here is to understand the backreaction of the metric due to the semiclassical energy fluxes inside an evaporating charged black hole. Recently an analytical approximation was developed for this semiclassical metric, with emphasis on the near-IH region – which we refer to as region 3. In this work, we sought to numerically validate this analytical approximation. To this end we developed a numerical code which evolves the semiclassical Einstein equation, from near the EH up to the very neighborhood of the IH.
We began with analyzing the Einstein equations and the Maxwell equations under spherical symmetry – from which we derived a convenient form of the evolution and constraint Einstein’s equations for the two metric variables $R(u,v)$ and $S(u,v)$ in double-null coordinates $u$,$v$. A finite-difference numerical code was developed to evolve the above mentioned evolution equations for $R$ and $S$ on a discrete grid. This code was constructed to be of second-order accuracy in the discretization parameter $du=dv$.
The dimensionless evaporation rate parameter is $\varepsilon_{0}$ ($\propto\frac{\hbar}{M^{2}}$). It is assumed to be $\ll 1$, corresponding to a macroscopic black hole. For astrophysical black holes, $\varepsilon_{0}$ is typically of order $10^{-82}$ or smaller, but in our numerical simulation due to obvious technical constraints we took it to be $\varepsilon_{0}=10^{-3}$ (if we wish to significantly evaporate the black hole within the numerical simulation, the grid size has to be of order $\frac{1}{\varepsilon_{0}}$ – which is totally infeasible for $\varepsilon_{0}\sim 10^{-82}$).
We checked the truncation error arising from the finite-$du$ grid size and found it to scale as $du^{2}$, as expected from our second-order numerical code. Overall the numerical error is small enough to allow us to accurately compare the numerical results with the analytical approximation.
In our numerical test of the analytical approximation, we focused primarily on the four derivatives $R_{,v}$$R_{,u}$$S_{,v}$$S_{,u}$. The numerical results confirmed the accuracy of the analytical approximation: we checked that the deviations between the numerical result and the analytical approximation for $S_{,u}$, $S_{,v}$ scale as $\varepsilon_{0}$, as expected. The deviations for $R_{,u}$$R_{,v}$ scale as $\varepsilon_{0}^{2}$, because $R_{,v}$ and $R_{,u}$ themselves are of order $\varepsilon_{0}$ (unlike $S_{,v}$ and $S_{,u}$ which scale as $(\varepsilon_{0})^{0}$).
As mentioned above, the analytical approximation for the metric inside the evaporating black hole is based on four main input functions (to be obtained from the semiclassical theory): $M(v)$ and $Q^{v}(v)$ at the event horizon, and $\widetilde{T}_{uu}^{(-)}(v),$$\text{$\widetilde{T}_{vv}^{(-)}$}(v)$ at the inner horizon. In order to run the simulation, we have to assume a specific form for the functions $T_{uv}^{(SC)}(u,v)$ and $T_{\theta\theta}^{(SC)}(u,v)$. (These functions, too, are to be determined from semiclassical gravity.) We found it more convenient to express these two stress-energy components via the source terms $Z_{R}$, $Z_{S}$. These source terms do affect the evolving flux components $T_{vv}$ and $T_{uu}$ – however they do not directly affect the evolving metric near the IH (provided that $\widetilde{T}_{vv}^{(-)}$ and $\widetilde{T}_{uu}^{(-)}$ have been specified). One of our goals was to directly test this independence. We verified this by running simulations with three different configurations of sources $Z_{R}$ and $Z_{S}$ which all lead to the same $\widetilde{T}_{vv}^{(-)}$, $\widetilde{T}_{uu}^{(-)}$. We found that all three source configurations produced the same results for the evolving metric in region 3 (up to $O(\varepsilon_{0})$ corrections) – as predicted by the analytical approximation.
In the next stage we run the code with three additional different configurations of the sources which lead to different $\widetilde{T}_{vv}^{(-)}$ and $\widetilde{T}_{uu}^{(-)}$ values, and have shown that the analytical approximation again gives good approximation for the metric variables in all these cases.
In one of the cases tested, $\widetilde{T}_{vv}^{(-)}$ and $\widetilde{T}_{uu}^{(-)}$ turned out to be positive. The analytical approximation then predicts that $R_{,v}$ and $R_{,u}$ should be negative throughout region 3. Therefore $R$ decreases to zero and a spacelike $R=0$ singularity should form. The numerical simulation confirmed this prediction. We presented the spacetime diagram in this case (see Figure 11.15) illustrating the formation of an $R=0$ singularity.
In conclusion, we numerically tested the validity of the analytical approximation for the backreaction metric inside an evaporating charged black hole in various situations. All these numerical simulations confirmed the validity of this analytical approximation.
Throughout this research we took the RSET to be of the form obtained by assuming RN background metric (in Unruh state). This was motivated by the expectation (or hope) that this RN-based RSET should well-approximate the RSET of the self-consistent solution of the semiclassical Einstein equation. It is not yet known if this expectation is realized or not. In any case, the backreaction metric constructed here nevertheless represents the result of the first stage in an iteration scheme aimed to construct the mentioned self-consistent solution. A natural future research direction would be to calculate the RSET in this backreaction metric, and in particular to see how similar or different is it from the RSET considered in this work.
12.1 Acknoledgments
I would like to thank my advisor Amos Ori. I would also like to thank Noa Zilberman. The financial help of the Technion is greatly appreciated.
13 Appendix A - irregularity of “naive $g_{1}$ metric”
We shall show here that the geometry given by the “naive metric”,
$ds^{2}=\left(1-\frac{2M(v)}{r}+\frac{Q(v)^{2}}{r^{2}}\right)du\;dv+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)$ (13.1)
is inherently irregular at the event horizon.
First of all, the determinant of the metric Eq. (13.1) is zero at $r=r_{+}(v)$ because
$1-\frac{2M(v)}{r_{+}(v)}+\frac{Q(v)^{2}}{r_{+}(v)^{2}}=0$ (13.2)
there. Unlike the standard Reissner–Nordström geometry in double null Eddington coordinates, in our case (where $M$ and $Q$ depend on v) this irregularity of the metric cannot be fixed by Kruskalization, as we shall now show.
Start by expanding the metric around $r_{+}$ (at fixed $v$):
$\left.\frac{\partial}{\partial r}\left(1-\frac{2M(v)}{r}+\frac{Q(v)^{2}}{r^{2}}\right)\right|_{r=r+}=\frac{2M}{r_{+}^{2}}-\frac{2Q^{2}}{r_{+}^{3}}=2\kappa_{+}^{v}(v)$ (13.3)
so around $r_{+}$the metric is approximated (at leading order in $r-r_{+}(v)$) by:
$ds^{2}=2\kappa_{+}^{v}(v)\left(r-r_{+}(v)\right)du\;dv+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)$ (13.4)
Around $r_{+},$along a $v=const$ line, $r_{*}$is approximated by (using Eq. (6.5)):
$\frac{u+v}{2}\equiv r_{*}=\frac{1}{2\kappa_{+}^{v}(v)}\log\left|r-r_{+}\right|+const(v)$ (13.5)
Solving for $r-r_{+}$:
$r-r_{+}=e^{2\kappa_{+}^{v}(v)r_{*}}const(v)=e^{\kappa_{+}^{v}(v)v+\kappa_{+}^{v}(v)u}const(v)$ (13.6)
Mark $const(v)\equiv h_{0}(v)$:
$r-r_{+}=e^{\kappa_{+}^{v}(v)v+\kappa_{+}^{v}(v)u}h_{0}(v)$ (13.7)
So $g_{uv}$ is equal to:
$g_{uv}=\kappa_{+}^{v}(v)\left(r-r_{+}\right)=\kappa_{+}^{v}(v)e^{\kappa_{+}^{v}(v)v+\kappa_{+}^{v}(v)u}h_{0}(v)$ (13.8)
and the overall metric is:
$ds^{2}=2\kappa_{+}^{v}(v)h_{0}(v)\;e^{\kappa_{+}^{v}(v)v}dv\;e^{\kappa_{+}^{v}(v)u}du+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)$ (13.9)
The issue will arise from the factor $e^{\kappa_{+}^{v}(v)u}$, which vanishes at the EH (where $u\rightarrow$-$\infty$). We shall therefore have to transform $u$ into a corresponding Kruskal-like coordinate $U$. But first we’ll transform $v$ to $V$ Kruskal coordinate for convenience. To this end, we need to absorb the exponent $e^{\kappa_{+}^{v}(v)v}$ into the definition of Kruskal coordinate, that is, we want to have:
$dV=e^{\kappa_{+}^{v}(v)v}dv$ (13.10)
Namely, $\frac{dV}{dv}=e^{\kappa_{+}^{v}(v)v}$, and by integrating we get:
$V(v)=\int e^{\kappa_{+}^{v}(v)v}dv$. We get the line element:
$ds^{2}=2h_{0}(v)\;\kappa_{+}^{v}(v)e^{\kappa_{+}^{v}(v)u}du\;dV+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)$ (13.11)
Note that $\kappa_{+}^{v}(v)$ is a function of $v$ and therefore it is a function of $V$.
We’ll now build the Kruskal coordinate $U$. $\frac{\partial U}{\partial u}$ needs to absorb the problematic term $e^{\kappa_{+}^{v}(v)u}$ in the metric. That means that $\frac{\partial U}{\partial u}=C(v)e^{\kappa_{+}^{v}(v)u}$ for some coefficient $C(v)$. From this we get by integration that $U=\frac{C(v)}{\kappa_{+}^{v}(v)}e^{\kappa_{+}^{v}(v)u}+Const(v)$. It turns out that both coefficients $C(v)$ and $Const(v)$ do not affect the regularity, so we choose
$U\equiv e^{\kappa_{+}(v)u}$ (13.12)
We get that:
$\frac{\partial U}{\partial u}=\kappa_{+}^{v}(v)e^{\kappa_{+}^{v}(v)u}$ (13.13)
$u$ satisfies:
$u=\frac{\log U}{\kappa_{+}^{v}(v)}$ (13.14)
[…]
$\frac{\partial u}{\partial V}=\frac{-1}{\kappa_{+}^{v}(v)^{2}}\;\frac{d}{dV}\left(\kappa_{+}^{v}(v)\right)\log U=h_{1}(V)\log U$ (13.15)
where we have defined
$h_{1}(V)\equiv\frac{-1}{\kappa_{+}^{v}(v)^{2}}\;\frac{d}{dV}\left(\kappa_{+}^{v}(v)\right)$ (13.16)
We calculate $g_{VV}$:
$\widehat{g}_{VV}$ $=g_{uV}\frac{\partial u}{\partial V}\frac{\partial V}{\partial V}=\kappa_{+}^{v}(v)h_{0}(v)Uh_{1}(V)\log U$ (13.17)
$=\kappa_{+}^{v}(v)h_{0}(v)h_{1}(V)U\log U=h_{2}(v)U\log U$ (13.18)
where we have defined $h_{2}(v)$ as:
$h_{2}(v)\equiv\kappa_{+}^{v}(v)h_{0}(v)h_{1}(V)$ (13.19)
Take the derivative of $\widehat{g}_{VV}$ with respect to $U$:
$\widehat{g}_{VV,U}=h_{2}(v)\left(\log U+1\right)$ (13.20)
Note that as $u\rightarrow$-$\infty$ at the event horizon, $U\rightarrow 0$, $\log U\rightarrow$-$\infty$, so $\widehat{g}_{VV,U}\rightarrow$-$\infty$. The $UV$ component of the new metric is:
$\widehat{g}_{UV}=g_{uV}\frac{\partial u}{\partial U}\frac{\partial V}{\partial V}=\kappa_{+}^{v}(v)h_{0}(v)U\frac{1}{U\kappa_{+}^{v}(v)}=h_{0}(v)$ (13.21)
Now the metric is:
$ds^{2}=h_{2}(V)U\log UdV^{2}+2h_{0}(v)dU\;dV+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)$ (13.22)
We see that while the metric is continuous and non-singular (namely $\det\widehat{g}\neq 0$) at the EH, the derivative of the metric diverges there.
To see that this divergence reflects a true geometrical singularity and not just coordinate singularity, we look at the scalar curvature of the metric. Using the Ricci package we get:
$R_{curvature}$ $=\frac{h_{2}(v)}{Uh_{0}(v)^{2}}+\frac{4h_{2}(v)r\;r_{,U}\log U}{r^{2}h_{0}(v)^{2}}$ (13.23)
$+\frac{1}{r^{2}h_{0}(v)^{2}}\left(2h_{0}(v)^{2}+4h_{2}(v)r\;r_{,U}-4h_{0}(v)r_{,V}r_{,U}+\right)$ (13.24)
$+\frac{1}{r^{2}h_{0}(v)^{2}}\left(2Uh_{2}(v)\kappa_{+}(v)r_{,U}^{2}\log U-8h_{0}(v)r\;r_{,UV}+4Uh_{2}(v)r\;r_{,UU}\log U\right)$ (13.25)
We wish to show that this expression diverges as $U\rightarrow 0$ at the EH. To do that, we calculate from Eq. (13.7) (the approximation of $r-r_{+}$ near the EH) the derivatives of $r$:
$r-r_{+}=e^{\kappa_{+}^{v}(v)v+\kappa_{+}^{v}(v)u}h_{0}(v)\equiv Uh_{3}(v)$ (13.26)
where we defined $h_{3}$ as:
$h_{3}(v)\equiv h_{0}(v)e^{\kappa_{+}^{v}(v)v}$ (13.27)
Therefore, the derivative $r_{,U}$ is equal to:
$r_{,U}=h_{3}(v)$ (13.28)
And the derivative $r_{,V}$ is equal to:
$r_{,V}=Uh_{3}^{\prime}(v)$ (13.29)
The second derivatives are ($r_{,VV}$ is not needed for this analysis):
$r_{,UV}=h_{3}^{\prime}(v),\hfill r_{,UU}=0$ (13.30)
Substituting these derivatives into the expression for the scalar curvature $R$ we obtain:
$R_{curvature}$ $=\frac{h_{2}(v)}{Uh_{0}(v)^{2}}+\frac{4h_{2}(v)rh_{3}(v)\log U}{r^{2}h_{0}(v)^{2}}$ (13.31)
$+\frac{1}{r^{2}h_{0}(v)^{2}}\left(2h_{0}(v)^{2}+4h_{2}(v)rh_{3}(v)-4h_{0}(v)Uh_{3}^{\prime}(v)h_{3}(v)\right)$ (13.32)
$+\frac{1}{r^{2}h_{0}(v)^{2}}\left(2Uh_{2}(v)\kappa_{+}(v)h_{3}(v)^{2}\log U-8h_{0}(v)rh_{3}^{\prime}(v)\right)$ (13.33)
As $U\rightarrow 0$, the first term diverges like $\frac{1}{U}$, the second term diverges weaker as $\log U$, and the rest is finite. We conclude that the scalar curvature $R_{curvature}$ diverges like $\frac{1}{U}$ at the EH, implying there is an inherent curvature singularity there.
14 Appendix B - determination of the initial values from $R_{,u}$, $S_{,u}$
We know the geometry in the near EH region 1, it is given by the $g_{1}$ metric Eq. (6.29,6.30). We use this metric to construct the initial conditions for the numerically evolving metric in region 2 (and subsequently region 3). In principle, this requires the initial values of the unknowns $R$,$S$ and their time derivatives on a hypersurface of constant $T\equiv\frac{u+v}{2}$, namely on $T=T_{0}$. We denote $Y\equiv\frac{v-u}{2}$. Thus the standard initial value problem would in principle require the four functions $R$, $R_{,T}$, $S$, $S_{,T}$ to produce a unique solution (where the partial derivatives are taken while keeping $Y$ fixed). Because $R_{,u}=\frac{1}{2}R_{,T}-\frac{1}{2}R_{,Y}$, and $R_{,Y}$ is in principle known along the initial hypersurface (from the initial function $R(Y)$), we can also use $R_{,u}$ for the initial values instead of $R_{,T}$. We prefer using $R_{,u}$ instead of $R_{,T}$ because it’s easier to calculate, as $M(v)$ is constant along lines of constant $v.$
In practice, our numerical algorithm requires two adjacent rows of initial values of $R$ and $S$ (but not of $R_{,T}$ and $S_{,T}$), namely the row $T=T_{0}$ and the subsequent row. We show in Figure 14.1 the two rows of initial values, $T=T_{0}$ in blue and the next one in red, which are required for evolution into the black points. We are given $R_{,u}$ and $S_{,u}$ at $T=T_{0}$ directly from the initial values of $g_{1}$ metric, and need to translate it into values of $R$ and $S$ on the red row.
Figure 14.1: Initial values rows
Figure 14.2: Initial values points
We proceed to calculate $R$ at the red point $(i,j)$ as shown in Figure 14.2 by calculating $R_{,u}$ at the green point $(i,j-\nicefrac{{1}}{{2}})$, and use that, and the value of $R$ at the blue point (i,j1)(i,j-1), to obtain $R$ at the red grid point $(i,j)$ (see Eq. (14.3) below). To obtain $R_{,u}(i,j-\nicefrac{{1}}{{2}})$, we first calculate $R_{,u}$ at the yellow point $R(i-\nicefrac{{1}}{{2}},j-\nicefrac{{1}}{{2}})$ by averaging:
$R_{,u}(i-\nicefrac{{1}}{{2}},j-\nicefrac{{1}}{{2}})=\frac{1}{2}\left(R_{,u}(i-1,j)+R_{,u}(i,j-1)\right)$ (14.1)
We know $R_{,u}(i-1,j)$ and $R_{,u}(i,j-1)$ because they are given by $R_{,u}$ in the metric $g_{1}$. Then we calculate $R_{,u}$ at the green point $(i,j-\nicefrac{{1}}{{2}})$ by using $R_{,uv}:$
$R_{,u}(i,j-\nicefrac{{1}}{{2}})=R_{,u}(i-\nicefrac{{1}}{{2}},j-\nicefrac{{1}}{{2}})+\frac{1}{2}dvR_{,uv}(i-\nicefrac{{1}}{{2}},j-\nicefrac{{1}}{{2}})$ (14.2)
in which we get $R_{,uv}$ by using the evolution equations. Finally, we can calculate $R(i,j)$ from $R_{,u}$$(i,j-\nicefrac{{1}}{{2}})$ and R(i,j1)R(i,j-1):
$R(i,j)=R(i,j-1)+du\;R_{,u}(i,j-\nicefrac{{1}}{{2}})$ (14.3)
To summarize, we got $R(i,j)$ at the new red row from $R$ and $R_{,u}$ (and $R_{,uv}$) at the previous blue row:
$R(i,j)=R(i,j-1)+du\left(\frac{1}{2}\left(R_{,u}(i-1,j)+R_{,u}(i,j-1)\right)+\frac{1}{2}dvR_{,uv}(i-\nicefrac{{1}}{{2}},j-\nicefrac{{1}}{{2}})\right)$ (14.4)
The same procedure is applied to obtain $S$ at the new red row, using $S_{,u}$ (and $S_{,uv}$). This procedure eventually provides $R$ and $S$ (in the whole numerical grid) in the required precision, which is second order in $du$.